### v1

# Unusual switch from low-temperature T-quadratic resistivity in the underdoped pseudogap phase of cuprate superconductors to low-temperature T-linear resistivity in the overdoped strange-metal phase

Abstract
The transport experiments demonstrate unambiguously a dramatic switch from the low-temperature linear in temperature (T-linear) resistivity in the overdoped strange-metal phase to the low-temperature quadratic in temperature (T-quadratic) resistivity in the underdoped pseudogap phase of cuprate superconductors, however, a consensus on the origin of this unusual switch is still lacking. Here the low-temperature resistivity in the underdoped pseudogap phase of cuprate superconductors is investigated using the Boltzmann transport equation. The low-temperature resistivity originates from the electron umklapp scattering mediated by the spin excitation. However, the dominant contribution to the low-temperature resistivity mainly comes from the antinodal umklapp scattering. In particular, a low temperature $T_{\rm scale}$ scales with $\Delta^{2}_{p}$ in the underdoped regime due to the opening of a momentum dependent spin pseudogap, where $\Delta_{p}$ is the minimal umklapp vector at the antinode. Notably, this $T_{\rm scale}$ as a function of doping presents a similar behavior of the antinodal spin pseudogap crossover temperature, i.e., $T_{\rm scale}$ decreases with the increase of doping in the underdoped regime, and then is reduced to a very low temperature in the overdoped regime. In the underdoped regime, the resistivity is T-quadratic in the low temperatures below $T_{\rm scale}$ with the strength of the T-quadratic resistivity that weakens as the doping is raised. However, in the overdoped regime, the resistivity is T-linear in the low temperatures above $T_{\rm scale}$. The result in this paper together with the recent study on the electrical transport in the overdoped regime therefore show that the electron umklapp scattering from a spin excitation responsible for the low-temperature T-linear resistivity in the overdoped regime naturally produces the low-temperature T-quadratic resistivity in the underdoped regime resulting from the opening of a momentum dependent spin pseudogap. Concomitantly, the low-temperature T-linear resistivity in the overdoped strange-metal phase is dramatically switched to the low-temperature T-quadratic resistivity in the underdoped pseudogap phase.
pacs:
74.25.Fy, 74.25.Nf, 74.20.Mn, 74.72.-h
I Introduction
It has become very clear that cuprate superconductors are among the most complicated systems studied in condensed matter physics Bednorz86 ; Damascelli03 ; Campuzano04 ; Fink07 ; Deutscher05 ; Devereaux07 ; Fischer07 . The complications stems mainly from the fact that the parent compound of cuprate superconductors is a form of non-conductor called a Mott insulator Bednorz86 ; Damascelli03 ; Campuzano04 ; Fink07 ; Deutscher05 ; Devereaux07 ; Fischer07 . Superconductivity then is achieved by chemically introducing charge carriers to this Mott insulator, which leads to that the physical properties of cuprate superconductors significantly depend on the extent of doping, and the regimes have been classified into the underdoped, optimally doped, and overdoped, respectively. After intensive investigations over more than three decades, a substantial amount of reliable and reproducible data has been accumulated by using many probes Damascelli03 ; Campuzano04 ; Fink07 ; Deutscher05 ; Devereaux07 ; Fischer07 ; Hussey08 ; Kastner98 ; Timusk99 ; Hufner08 ; Vishik18 ; Fujita12 , which reveals that the most of the nonconventional and unique features of cuprate superconductors are observed in the normal-state above the superconducting (SC) transition temperature $T_{\rm c}$, where the deepest mysteries lie. In particular, below a pseudogap crossover temperature $T^{*}$, which can be well above $T_{\rm c}$ in the underdoped regime, the physical response can be well interpreted in terms of the formation of a pseudogap by which it means a large suppression of the electronic density of states at the electron Fermi surface (EFS) Damascelli03 ; Campuzano04 ; Fink07 ; Deutscher05 ; Devereaux07 ; Fischer07 ; Hussey08 ; Kastner98 ; Timusk99 ; Hufner08 ; Vishik18 ; Fujita12 . This is why in the underdoped regime, the phase above $T_{\rm c}$ but below $T^{*}$ is so-called as the pseudogap phase. The pseudogap in the underdoped regime was first discovered through Knight-shift of the magnetic susceptibility experiments Warren89 ; Alloul89 , where the pseudogap is some kind of the spin pseudogap, meaning a strong reduction in the magnetic susceptibility through the particle-hole correlation. Subsequently, the presence of the pseudogap was confirmed by a series of experimental measurements taken with a wide variety of techniques Walstedt90 ; Takigawa91 ; Loeser96 ; Norman98 ; Renner98 ; Kohsaka08 ; Puchkov96 . Moreover, these experiments also show clearly that the pseudogap exists in both the spin and charge channels Timusk99 ; Hufner08 ; Vishik18 ; Fujita12 ; Warren89 ; Alloul89 ; Walstedt90 ; Takigawa91 ; Loeser96 ; Norman98 ; Renner98 ; Kohsaka08 ; Puchkov96 . For example, angle-resolved photoemission spectroscopy (ARPES) and scanning tunneling spectroscopy measure the charge channel Loeser96 ; Norman98 , while nuclear magnetic resonance (NMR) and nuclear quadrupole resonance (NQR) detect the spin channelFujita12 ; Warren89 ; Alloul89 ; Walstedt90 ; Takigawa91 . On the other hand, in the overdoped regime, the normal-state is characterized by a number of the anomalous low-temperature properties Damascelli03 ; Campuzano04 ; Fink07 ; Deutscher05 ; Devereaux07 ; Fischer07 ; Hussey08 ; Kastner98 ; Timusk99 ; Hufner08 ; Vishik18 ; Fujita12 in the sense that they do not fit in with the conventional Fermi-liquid theory Schrieffer64 ; Abrikosov88 ; Mahan81 . This has led to the normal-state in the overdoped regime being refereed to as the strange-metal phaseKeimer15 .
Among the notable characteristics in the normal-state of cuprate superconductors, the most iconic feature has to be the low-temperature resistivity Hussey08 ; Kastner98 ; Timusk99 ; Hufner08 . In the overdoped regime, a series of transport measurements revealed a linear in temperature (T-linear) resistivity at low temperaturesKeimer15 ; Hussey23 ; Legros19 ; Ayres21 ; Grisso21 , which is in a striking contrast to the conventional Fermi-liquid behaviour found in the conventional superconductorsSchrieffer64 ; Abrikosov88 ; Mahan81 . However, in the underdoped regime, the early transport measurements indicated that the opening of the pseudogap in the spin excitation spectrum leads to a striking deviation from the low-temperature T-linear behaviour of the resistivityGurvitch87 . In particular, the subsequent transport experiments observed a reduced resistivity below $T^{*}$, near the temperature where the spin pseudogap, as seen by NMR Knight-shift and spin-relaxation rate, openedBucher93 ; Ito93 ; Nakano94 ; Ando01 . Thus the low-temperature resistivity anomaly coincides roughly with $T^{*}$ and these experimental results suggest that the decrease in the low-temperature below $T^{*}$ is caused by the reduced electron scattering by the spin excitations resulting from the opening the spin pseudogapBucher93 ; Ito93 ; Nakano94 ; Ando01 . Lately, the low-temperature resistivity was confirmed experimentally in a quadratic temperature dependenceAndo04a ; Ando04 ; Lee05 ; Proust08 , as would be expected from the conventional Fermi-liquid theory, for a wide doping range in the underdoped regime. More specifically, the recent transport experiments demonstrate unambiguously a clear and dramatic switch from the low-temperature T-linear resistivity in the overdoped regime to the purely quadratic in temperature (T-quadratic) resistivity in the underdoped regime Cooper09 ; Hussey11 ; Barisic13 ; Pelc20 , where the strength of the T-quadratic resistivity (then the T-quadratic resistivity coefficient) decreases with the increase of doping. In spite of the considerable variation in crystal structures and impurity-scattering effects among different families of cuprate superconductors, these experimental results Gurvitch87 ; Bucher93 ; Ito93 ; Nakano94 ; Ando01 ; Ando04a ; Ando04 ; Lee05 ; Proust08 ; Cooper09 ; Hussey11 ; Barisic13 ; Pelc20 also show clearly that as the low-temperature T-linear resistivity in the overdoped regimeLegros19 ; Ayres21 ; Grisso21 , the low-temperature T-quadratic resistivity also is a universal feature in the underdoped regime. In this case, some crucial questions are raised: (i) why the low-temperature resistivity exhibits a dramatic switch from the T-linear resistivity in the overdoped regime to the T-quadratic resistivity in the underdoped regime? (ii) whether the pseudogap is correlated with this dramatic switch or not? (iii) is there a common electron scattering mediated by the same bosonic mode that is responsible for both the low-temperature T-quadratic resistivity in the underdoped regime and low-temperature T-linear resistivity in the overdoped regime?
The theoretical explanation for this unusual switch from the low-temperature T-linear resistivity in the overdoped regime to the low-temperature T-quadratic resistivity in the underdoped regime along with the underlying scattering mechanism is clearly a big challenge. For a understanding of the nature of the low-temperature T-linear resistivity in the overdoped regime, several scenarios have been proposed Varma89 ; Varma16 ; Varma20 ; Damle97 ; Sachdev11 ; Zaanen04 ; Luca07 ; Zaanen19 ; Hartnoll22 ; Lee21 . However, these scenarios are just as diverse as the mechanism of superconductivity. In the marginal Fermi-liquid phenomenologyVarma89 ; Varma16 ; Varma20 , a single T-linear scattering rate is posited to account for the T-linear resistivity. In particular, the T-linear resistivity was interpreted as a consequence of the scale invariant physics near to the quantum critical point (QCP) Varma20 ; Damle97 ; Sachdev11 . With the close relation to the physics of QCP, the T-linear resistivity has been interpreted in terms of the Planckian dissipationZaanen04 ; Luca07 ; Zaanen19 ; Hartnoll22 . Moreover, it has been shown recently that the low-temperature resistivity originates from the umklapp scattering between electrons by the exchange of a critical boson propagatorLee21 , where the umklapp scattering rate is T-linear in low temperatures near the umklapp point, which induces a low-temperature T-linear resistivity. These studiesLee21 together with many othersHonerkamp01 ; Hartnoll12 thus suggest that the electron umklapp scattering is the origin of the low-temperature T-linear resistivity. On the other hand, it has been argued that if the electron scattering responsible for the low-temperature T-linear resistivity in the overdoped regime involves the electron scattering on the spin excitations in the underdoped regime Puchkov96 , then the spin pseudogap seen in NMR below $T^{*}$ would naturally account for the deviation from the low-temperature T-linear behaviour of the resistivityBucher93 ; Ito93 ; Nakano94 ; Ando01 . Moreover, the mechanism for the resistivity in the underdoped regime has been proposedHussey03 ; Rice17 , where the electron umklapp scattering above $T^{*}$ leads to the T-linear resistivity Rice17 . As the temperatures fall below $T^{*}$, the pseudogap opens, and then restricts the available umklapp scattering channels, leading to a crossover between the T-linear resistivity at the temperature above $T^{*}$ and the traditional T-quadratic form of the scattering from the Fermi pockets at the temperature below $T^{*}$. However, up to now, the microscopic origin of the low-temperature T-quadratic resistivity in the underdoped regime and its connection with the pseudogap are still much debated, and no explanation of the unusual switch from the low-temperature T-linear resistivity in the overdoped regime to the low-temperature T-quadratic resistivity in the underdoped regime has been made so far starting from a microscopic theory.
In the very recent studyMa23 , we have investigated the nature of the low-temperature T-linear resistivity in the overdoped strange-metal phase. In our scenario, the scattering rate is arisen from the umklapp scattering between electrons by the exchange of the effective spin propagator, where the dominant contribution to the resistivity mainly comes from the antinodal umklapp scattering. This umklapp scattering rate scales linearly with temperature in the low temperatures, which then naturally generates a low-temperature T-linear resistivity. In this paper, we study the low-temperature electrical transport in the underdoped pseudogap phase along with this line, where in a striking difference to the case in the overdoped regime, the spin excitation energy dispersion is anisotropically renormalized due to the opening of a momentum dependent spin pseudogap. In particular, the density of the spin excitations at around the antinodal is heavily reduced by the antinodal spin pseudogap. In this case, a low temperature $T_{\rm scale}$ scales with $\Delta^{2}_{p}$, where $\Delta_{p}$ is the minimal umklapp vector at the antinode. Notably, this low-temperature scale $T_{\rm scale}$ as a function of doping presents a similar behavior of the antinodal spin pseudogap crossover temperature, i.e., $T_{\rm scale}$ is dropped down with the enhancement of doping in the underdoped regime, and then is reduced to a very low temperature in the overdoped regime. In the underdoped regime, the resistivity is T-quadratic at the low temperatures below $T_{\rm scale}$, where the T-quadratic resistivity strength weakens as the doping is raised. However, in the overdoped regime, the resistivity is instead T-linear in the low temperatures above $T_{\rm scale}$. The present results combined with the recent findingsMa23 on the electrical transport in the overdoped regime therefore reveal that (i) the electron umklapp scattering from a spin excitation associated with the antinodes leads to the T-linear resistivity in the weak coupling overdoped regime; (ii) as this electron umklapp scattering flows to the strong coupling underdoped regime, the opening of the momentum dependent spin pseudogap lowers the density of the spin excitations at around the antinodal region in the response to the umklapp scattering, which reduces the strength of the intense umklapp scattering from the electronic states into the antinodal region, and therefore leads to the low-temperature T-quadratic form of the umklapp scattering rate. Consequently, the low-temperature T-linear resistivity in the overdoped strange-metal phase is dramatically switched to the low-temperature T-quadratic resistivity in the underdoped pseudogap phase.
This paper is organized as follows. In the next section, we first provide a brief summary for the derivation of the full spin and electron propagators, where (i) the antinodal spin pseudogap decreases with the increase of doping in the underdoped regime, and then it abruptly vanishes at around the optimal doping; (ii) the electronic density of state at around the antinodal region is gapped out by the normal-state pseudogap, and then the closed EFS contour is truncated to a set of four disconnected Fermi arcs centered at around the nodal region. Following this full spin propagator and the reconstructed EFS, the scattering rate originated from the umklapp scattering between electrons by the exchange of the full effective spin propagator is derived using the Boltzmann transport equation. The discussions of the quantitative characteristics of the low-temperature resistivity in the underdoped pseudogap phase are given in Section III, where we show that the scattering rate in the whole doping regime is predominantly governed by the antinodal umklapp scattering. In Section IV, we close the presentation with a summary and discussion of the main results.
II Methodology
II.1 Low-energy effective $t$-$J$ model
The common element in the layered crystal structure of cuprate superconductors is the square-lattice copper-oxide layerBednorz86 , and then the nonconventional features of cuprate superconductors mainly come from the strongly correlated motion of the electrons in the copper-oxide layerCooper94 ; Nakamura93 ; Hou94 ; Takenaka94 . Quickly after the discovery of superconductivity in cuprate superconductors, Anderson Anderson87 recognized that the essential physics of the doped copper-oxide layer can be modeled with the low-energy effective $t$-$J$ model,
$H$ $=$ $-t\sum_{\langle l\hat{\eta}\rangle\sigma}C^{\dagger}_{l\sigma}C_{l+\hat{\eta}\sigma}+t^{\prime}\sum_{\langle l\hat{\tau}\rangle\sigma}C^{\dagger}_{l\sigma}C_{l+\hat{\tau}\sigma}+\mu\sum_{l\sigma}C^{\dagger}_{l\sigma}C_{l\sigma}$ (1)
$+$ $J\sum_{\langle l\hat{\eta}\rangle}{\bf S}_{l}\cdot{\bf S}_{l+\hat{\eta}},~{}~{}~{}~{}$
where $C^{\dagger}_{l\sigma}$ and $C_{l\sigma}$ are electron operators that respectively create and annihilate an electron on site $l$ with spin $\sigma$, ${\bf S}_{l}$ is the spin operator of the electron with its components $S_{l}^{x}$, $S_{l}^{y}$, and $S_{l}^{z}$, and $\mu$ is the chemical potential. This $t$-$J$ model (1) describes a competition between the kinetic energy and magnetic energy, and the kinetic energy includes the nearest-neighbor (NN) hopping with the hopping integral $t$ and the next NN hopping with the hopping integral $t^{\prime}$, while the magnetic energy is described by an antiferromagnetic (AF) Heisenberg term with the NN coupling $J$. $\langle l\hat{\eta}\rangle$ ($\langle l\hat{\tau}\rangle$) indicates that $l$ runs over all sites, and for each $l$, over its NN sites $\hat{\eta}$ (next NN sites $\hat{\tau}$). The parameters in this paper are chosen as $t/J=2.5$ and $t^{\prime}/t=0.3$ as the recent discussions Ma23 . Moreover, we use a notation in which the magnetic coupling $J$ and the lattice constant of the square lattice are set to the energy and length units, respectively.
The $t$-$J$ model (1) is supplemented by a crucial on-site local constraint $\sum_{\sigma}C^{\dagger}_{l\sigma}C_{l\sigma}\leq 1$ that the double occupancy of a site by two electrons of opposite spins is not allowed Anderson87 , while the strong electron correlation in the system manifests itself by this no double electron occupancy local constraint Yu92 ; Lee06 ; Edegger07 ; Spalek22 ; Zhang93 . However, the no double electron occupancy also leads to the difficulty for the studying of the $t$-$J$ model (1). For a proper treatment of this no double electron occupancy local constraint, the fermion-spin transformationFeng0494 ; Feng15 has been established, where the physics of no double electron occupancy is taken into account by the representation of the constrained electron as,
$C_{l\uparrow}=h^{\dagger}_{l\uparrow}S^{-}_{l},~{}~{}~{}~{}C_{l\downarrow}=h^{\dagger}_{l\downarrow}S^{+}_{l},$ (2)
where the $U(1)$ gauge invariant spin-raising (spin-lowering) operator $S^{+}_{l}$ ($S^{-}_{l}$) carries spin index of the constrained electron, and therefore the collective mode from this spin degree of freedom of the constrained electron is interpreted as the spin excitation responsible for the spin dynamics of the system, and the $U(1)$ gauge invariant spinful fermion operator $h^{\dagger}_{l\sigma}=e^{i\Phi_{l\sigma}}h^{\dagger}_{l}$ ($h_{l\sigma}=e^{-i\Phi_{l\sigma}}h_{l}$) creates (annihilates) a charge carrier on site $l$, and therefore represents the charge degree of freedom of the constrained electron together with some effects of spin configuration rearrangements due to the presence of the doped charge carrier itself, while the charge carrier and localized spin recombine to form the physical electron responsible for the electronic properties. Following this fermion-spin transformation (2), we can rewrite the $t$-$J$ model (1) as,
$H$ $=$ $\sum_{\langle l\hat{\eta}\rangle}t(h^{\dagger}_{l+\hat{\eta}\uparrow}h_{l\uparrow}S^{+}_{l}S^{-}_{l+\hat{\eta}}+h^{\dagger}_{l+\hat{\eta}\downarrow}h_{l\downarrow}S^{-}_{l}S^{+}_{l+\hat{\eta}})$ (3)
-- $\sum_{\langle l\hat{\tau}\rangle}t^{\prime}(h^{\dagger}_{l+\hat{\tau}\uparrow}h_{l\uparrow}S^{+}_{l}S^{-}_{l+\hat{\tau}}+h^{\dagger}_{l+\hat{\tau}\downarrow}h_{l\downarrow}S^{-}_{l}S^{+}_{l+\hat{\tau}})$
-- $\mu_{\rm h}\sum_{l\sigma}h^{\dagger}_{l\sigma}h_{l\sigma}+J_{\rm eff}\sum_{\langle l\hat{\eta}\rangle}{\bf S}_{l}\cdot{\bf S}_{l+\hat{\eta}},$
with $J_{{\rm eff}}=(1-\delta)^{2}J$, the charge-carrier doping concentration $\delta=\langle h^{\dagger}_{l\sigma}h_{l\sigma}\rangle=\langle h^{\dagger}_{l}h_{l}\rangle$, and the charge-carrier chemical potential $\mu_{\rm h}$. The above $t$-$J$ model (3) therefore describes a doped Mott insulator as a sparse density of the charge carriers moving in a background of an AF coupled square lattice of spins, while the motion of the electrons rearranges the spin configuration leading to the strong coupling between the charge and spin degrees of freedom of the constrained electron.
II.2 Pseudogap in charge and spin channels
In the past three decades, a series of experiments from NMR, NQR, and the inelastic neutron scattering measurements Fujita12 ; Birgeneau89 ; Fong95 ; Yamada98 ; Arai99 ; Bourges00 ; He01 ; Hayden04 ; Tranquada04 ; Bourges05 has provided an intrinsic connection between the electron pairing mechanism and spin excitations in cuprate superconductors, where a key question is whether the spin excitation can mediate electron pairing in analogy to the phonon-mediate pairing mechanism in the conventional superconductorsSchrieffer64 . Starting from the fermion-spin formulation of the $t$-$J$ model (3), the kinetic-energy-driven superconductivityFeng15 ; Feng0306 ; Feng12 ; Feng15a has been developed, where the d-wave charge-carrier pairing state is generated by the charge-carrier interaction directly from the kinetic energy of the $t$-$J$ model (3) by the exchange of the spin excitation. However, the d-wave electron pairs originated from this d-wave charge-carrier pairing state are due to the charge-spin recombinationFeng15a , and these electron pairs condensation reveals the d-wave SC-state. This kinetic-energy-driven SC mechanism is purely electronic without phonon, since the glue to hold the constrained electron pairs together is the spin excitation, the collective mode from the spin degree of freedom of the constrained electron itself. Moreover, this electron paring state mediated by the spin excitation in a way is in turn strongly influenced by the electron coherence, which leads to that the maximal $T_{\rm c}$ occurs at the optimal doping, and then decreases in both the underdoped and overdoped regimes. In the kinetic-energy-driven superconductivity, the self-consistent equations that are satisfied by the full charge-carrier diagonal and off-diagonal propagators in the charge-carrier pairing state have been evaluated in terms of the Eliashberg formalismEliashberg60 ; Scalapino66 , and then in the charge-carrier normal-state, these self-consistent equations are reduced asFeng15 ; Feng12 ,
$g({\bf k},\omega)=g^{(0)}({\bf k},\omega)+g^{(0)}({\bf k},\omega)\Sigma^{({\rm h})}_{\rm ph}({\bf k},\omega)g({\bf k},\omega),~{}~{}~{}~{}$ (4)
with the charge-carrier propagator of the $t$-$J$ model (3) in the mean-filed (MF) approximation $g^{(0)-1}({\bf k},\omega)=\omega-\xi_{\bf k}$, where $\xi_{\bf k}=4t\chi_{1}\gamma_{{\bf k}}-4t^{\prime}\chi_{2}\gamma_{{\bf k}}^{\prime}-\mu_{\rm h}$ is the MF charge-carrier energy dispersion, with $\gamma_{\bf k}=({\rm cos}k_{x}+{\rm cos}k_{y})/2$, $\gamma_{\bf k}^{\prime}={\rm cos}k_{x}{\rm cos}k_{y}$, and the spin correlation functions $\chi_{1}=\langle S^{+}_{l}S^{-}_{l+\hat{\eta}}\rangle$ and $\chi_{2}=\langle S_{l}^{+}S_{l+\hat{\tau}}^{-}\rangle$. From the above self-consistent equation (4), the full charge-carrier propagator can be expressed explicitly as,
$g({\bf k},\omega)={1\over\omega-\xi_{\bf k}-\Sigma^{({\rm h})}_{\rm ph}({\bf k},\omega)},$ (5)
where the charge-carrier normal self-energy is given by,
$\Sigma^{({\rm h})}_{\rm ph}({\bf k},i\omega_{n})$ $=$ ${1\over N}\sum_{\bf p}{1\over\beta}\sum_{ip_{m}}g({\bf p}+{\bf k},ip_{m}+i\omega_{n})$ (6)
$\times$ $P^{(0)}({\bf k},{\bf p},ip_{m}),~{}~{}~{}$
with the number of lattice sites N, the fermion and bosonic Matsubara frequencies $\omega_{n}$ and $p_{m}$, respectively, and the MF effective spin propagator,
$P^{(0)}({\bf k},{\bf p},\omega)={1\over N}\sum_{\bf q}\Lambda^{2}_{{\bf p}+{\bf q}+{\bf k}}\Pi({\bf p},{\bf q},\omega),$ (7)
where $\Lambda_{{\bf k}}=4t\gamma_{\bf k}-4t^{\prime}\gamma_{\bf k}^{\prime}$ is the bare vertex function, while the spin bubble $\Pi({\bf p},{\bf q},\omega)$ is a convolution of two MF spin propagators, and can be expressed as,
$\Pi({\bf p},{\bf q},ip_{m})={1\over\beta}\sum_{iq_{m}}D^{(0)}({\bf q},iq_{m})D^{(0)}({\bf q}+{\bf p},iq_{m}+ip_{m}),~{}~{}~{}$ (8)
with the bosonic Matsubara frequency $q_{m}$, and the MF spin propagator,
$D^{(0)}({\bf k},\omega)={B_{\bf k}\over\omega^{2}-\omega^{2}_{\bf k}}={B_{\bf k}\over 2\omega_{\bf k}}\left({1\over\omega-\omega_{\bf k}}-{1\over\omega+\omega_{\bf k}}\right),~{}~{}~{}~{}~{}$ (9)
where the MF spin excitation energy dispersion $\omega_{\bf k}$ and the corresponding weight function of the MF spin excitation spectrum $B_{\bf k}$ have been given explicitly in the previous worksFeng15 ; Feng0306 ; Feng12 .
Figure 1: (a) The charge-carrier pseudogap at temperature $T=0.002J$ and (b) the charge-carrier pseudogap crossover temperature as a function of doping taken from Ref. Feng12, .
With the help of the above spin propagator (9), the charge-carrier normal self-energy $\Sigma^{({\rm h})}_{\rm ph}({\bf k},\omega)$ in Eq. (6) has been derived asFeng15 ; Feng0306 ; Feng12 ,
$\Sigma^{({\rm h})}_{\rm ph}({\bf k},\omega)$ $=$ ${1\over N^{2}}\sum_{{\bf pp}^{\prime}\mu\nu}(-1)^{\nu+1}Z^{({\rm h})}_{\rm F}\Omega_{{\bf p}{\bf p}^{\prime}{\bf k}}$ (10)
$\times$ ${F^{({\rm h})}_{\mu\nu}({\bf p},{\bf p}^{\prime},{\bf k})\over\omega+(-1)^{\mu+1}\omega^{(\nu)}_{{\bf p}{\bf p}^{\prime}}-\bar{\xi}_{{\bf p}+{\bf k}}},~{}~{}~{}$
with $\mu~{}(\nu)=1,~{}2$, $\omega^{(\nu)}_{{\bf p}{\bf p}^{\prime}}=\omega_{{\bf p}+{\bf p}^{\prime}}-(-1)^{\nu}\omega_{\bf p^{\prime}}$, $\Omega_{{\bf p}{\bf p}^{\prime}{\bf k}}=\Lambda^{2}_{{\bf p}+{\bf p}^{\prime}+{\bf k}}B_{{\bf p}^{\prime}}B_{{\bf p}+{\bf p}^{\prime}}/[4\omega_{{\bf p}^{\prime}}\omega_{{\bf p}+{\bf p}^{\prime}}]$, $\bar{\xi_{{\bf k}}}=Z^{({\rm h})}_{\rm F}\xi_{{\bf k}}$, $F^{({\rm h})}_{\mu\nu}({\bf p},{\bf p}^{\prime},{\bf k})=n_{\rm F}[(-1)^{\mu+1}\bar{\xi}_{{\bf p}+{\bf k}}]n^{(\nu)}_{{\rm 1B}{{\bf p}{\bf p}^{\prime}}}+n^{(\nu)}_{{\rm 2B}{{\bf p}{\bf p}^{\prime}}}$, $n^{(\nu)}_{{\rm 1B}{{\bf p}{\bf p}^{\prime}}}=1+n_{\rm B}(\omega_{{\bf p}^{\prime}+{\bf p}})+n_{\rm B}[(-1)^{\nu+1}\omega_{\bf p^{\prime}}]$, $n^{(\nu)}_{{\rm 2B}{{\bf p}{\bf p}^{\prime}}}=n_{\rm B}(\omega_{{\bf p}^{\prime}+{\bf p}})n_{\rm B}[(-1)^{\nu+1}\omega_{\bf p^{\prime}}]$, the boson and fermion distribution functions $n_{\rm B}(\omega)$ and $n_{\rm F}(\omega)$, respectively, while the charge-carrier coherent weight $Z^{({\rm h})}_{\rm F}$ that has been obtained in the previous works Feng15 ; Feng0306 ; Feng12 . Moreover, as the same approach of the numerical calculation described in Ref. Ma23, , the calculation of the normal self-energy and the related quantities in this paper is also performed numerically on a $160\times 160$ lattice in momentum space.
In the framework of the kinetic-energy-driven superconductivity Feng15 ; Feng0306 ; Feng12 ; Feng15a , the charge-carrier pseudogap forms due to the strong coupling between charge and spin degrees of freedom of the constrained electron. To explore the nature of this charge-carrier pseudogap more clearly, we rewrite the above charge-carrier normal self-energy (10) as,
$\Sigma^{({\rm h})}_{\rm ph}(\bf k,\omega)$$\approx$ ${[\bar{\Delta}^{({\rm h})}_{\rm pg}({\bf k})]^{2}\over\omega-\xi_{0{\bf k}}},$ (11)
where $\xi_{0{\bf k}}=L^{({\rm h})}_{2}({\bf k})/L^{({\rm h})}_{1}({\bf k})$ is the energy spectrum of $\Sigma^{({\rm h})}_{\rm ph}(\bf k,\omega)$, and $\bar{\Delta}^{({\rm h})}_{\rm pg}({\bf k})=L^{({\rm h})}_{2}({\bf k})/\sqrt{L^{({\rm h})}_{1}({\bf k})}$ is refereed to as the momentum dependence of the charge-carrier pseudogap, since it plays a role of the suppression of the low-energy spectral weight of the charge-carrier excitation spectrum, with the functions $L^{({\rm h})}_{1}({\bf k})=-\Sigma^{({\rm h})}_{\rm pho}({\bf k},\omega=0)$ and $L^{({\rm h})}_{2}({\bf k})=-\Sigma^{({\rm h})}_{\rm ph}({\bf k},\omega=0)$, while $\Sigma^{({\rm h})}_{\rm pho}({\bf k},\omega=0)$ and the antisymmetric part $\Sigma^{({\rm h})}_{\rm pho}({\bf k},\omega)$ of the charge-carrier normal self-energy can be derived directly from the charge-carrier normal self-energy $\Sigma^{({\rm h})}_{\rm ph}({\bf k},\omega)$ in Eq. (10).
For a convenience in the following discussions, the above charge-carrier pseudogap Feng15 ; Feng12 $\bar{\Delta}^{({\rm h})}_{\rm pg}$ at temperature $T=0.002J$ as a function of doping is replotted in Fig. 1a, where the relatively large charge-carrier pseudogap appears in the underdoped regime, and then it weakens as the optimal doping is approached. However, a quite weak charge-carrier pseudogap is still present at around the optimal doping, but it disappears at the heavily overdoped region. Moreover, at a given doping concentration, this charge-carrier pseudogap is identified with a crossover with a charge-carrier pseudogap crossover temperature $T^{*}_{\rm h}$ rather than a phase transition. To see the evolution of this $T^{*}_{\rm h}$ with doping more clearly, $T^{*}_{\rm h}$ as a function of doping is also replotted in Fig. 1b, where in conformity with the doping dependence of the charge-carrier pseudogap in Fig. 1a, $T^{*}_{\rm h}$ is relatively high at the slight underdoping, and then it smoothly decreases with the increase of doping in the underdoped regime, eventually terminating at the heavily overdoped regionFeng12 ; Feng15 . More importantly, as the electron pairing state originated from the charge-carrier pairing state are due to the charge-spin recombination, the normal-state pseudogap state originated from the charge-carrier pseudogap state is also due to the charge-spin recombinationFeng15a , and then the anomalous properties associated with the formation of the normal-state pseudogap are explained in a natural way Damascelli03 ; Campuzano04 ; Fink07 ; Deutscher05 ; Devereaux07 ; Fischer07 ; Hussey08 ; Kastner98 ; Timusk99 ; Hufner08 ; Vishik18 ; Fujita12 . We will return to the discussion of the normal-state pseudogap towards next subsection II.3.
In the fermion-spin theoryFeng0494 ; Feng15 , the scattering of spins due to the charge-carrier fluctuation dominates the spin dynamics. In this case, the dynamical spin response of cuprate superconductors has been discussed from the SC-state to normal-stateKuang15 ; Yuan01 ; Feng98 , where the full spin propagator in the normal-state has been evaluated as,
$D({\bf k},\omega)$ $=$ ${1\over D^{(0)-1}({\bf k},\omega)-\Sigma^{({\rm s})}_{\rm ph}({\bf k},\omega)}$ (12)
$=$ ${B_{\bf k}\over\omega^{2}-\omega^{2}_{\bf k}-B_{\bf k}\Sigma^{({\rm s})}_{\rm ph}({\bf k},\omega)},~{}~{}~{}~{}~{}$
with the spin self-energy in the normal-state that is derived in terms of the collective charge-carrier mode in the particle-hole channel asKuang15 ,
$\Sigma^{({\rm s})}_{\rm ph}({\bf k},\omega)$ $=$ $-{1\over N^{2}}\sum_{\bf pq}\Omega^{({\rm s})}_{{\bf k}{\bf p}{\bf q}}{F^{\rm(s)}({\bf k},{\bf p},{\bf q})\over\omega^{2}-[\omega_{{\bf q}+{\bf k}}-(\bar{\xi}_{{\bf p}+{\bf q}}-\bar{\xi}_{{\bf p}})]^{2}},$
where the functions $\Omega^{({\rm s})}_{{\bf k}{\bf p}{\bf q}}$ and $F^{\rm(s)}({\bf k},{\bf p},{\bf q})$ are given by,
$\Omega^{({\rm s})}_{{\bf k}{\bf p}{\bf q}}$ $=$ ${B_{{\bf q}+{\bf k}}\over\omega_{{\bf q}+{\bf k}}}[Z^{({\rm h})}_{\rm F}]^{2}(\Lambda^{2}_{{\bf k}-{\bf p}}+\Lambda^{2}_{{\bf p}+{\bf q}+{\bf k}}),~{}~{}~{}~{}~{}~{}$ (14a)
$F^{\rm(s)}({\bf k},{\bf p},{\bf q})$ $=$ $[\omega_{{\bf q}+{\bf k}}-(\bar{\xi}_{{\bf p}+{\bf q}}-\bar{\xi}_{{\bf p}})]\{n_{\rm B}(\omega_{{\bf q}+{\bf k}})[n_{\rm F}(\bar{\xi}_{{\bf p}})$ (14b)
-- $n_{\rm F}(\bar{\xi}_{{\bf p}+{\bf q}})]-[1-n_{\rm F}(\bar{\xi}_{{\bf p}})]n_{\rm F}(\bar{\xi}_{{\bf p}+{\bf q}})\},~{}~{}~{}~{}~{}~{}~{}~{}$
respectively. For a better understanding of the unusual features of the spin pseudogap, the above spin self-energy (II.2) can be also rewritten as,
$\Sigma^{({\rm s})}_{\rm ph}({\bf k},\omega)\approx{B_{\bf k}[\bar{\Delta}^{({\rm s})}_{\rm pg}({\bf k})]^{2}\over\omega^{2}-\omega^{2}_{0{\bf k}}},$ (15)
where $\omega^{2}_{0{\bf k}}=L^{({\rm s})}_{1{\bf k}}/L^{({\rm s})}_{2{\bf k}}$ is the energy spectrum of $\Sigma^{({\rm s})}_{\rm ph}({\bf k},\omega)$, while $[\bar{\Delta}^{({\rm s})}_{\rm pg}({\bf k})]^{2}=[L^{({\rm s})}_{1{\bf k}}]^{2}/B_{\bf k}L^{({\rm s})}_{2{\bf k}}$ is identified as being a region of the spin self-energy in which $\bar{\Delta}^{({\rm s})}_{\rm pg}({\bf k})$ anisotropically reduces the density of the spin excitation, and in this sense, $\bar{\Delta}^{({\rm s})}_{\rm pg}({\bf k})$ is refereed to as the spin pseudogap. The functions $L^{({\rm s})}_{1{\bf k}}$ and $L^{({\rm s})}_{2{\bf k}}$ can be derived directly from the spin self-energy $\Sigma^{({\rm s})}_{\rm ph}({\bf k},\omega)$ in Eq. (II.2) as,
$L^{({\rm s})}_{1{\bf k}}$ $=$ $-{B_{\bf k}\over N^{2}}\sum_{\bf pq}\Omega^{({\rm s})}_{{\bf k}{\bf p}{\bf q}}{F^{\rm(s)}({\bf k},{\bf p},{\bf q})\over[\omega_{{\bf q}+{\bf k}}-(\bar{\xi}_{{\bf p}+{\bf q}}-\bar{\xi}_{{\bf p}})]^{2}},~{}~{}~{}~{}~{}$ (16a)
$L^{({\rm s})}_{2{\bf k}}$ $=$ $-{B_{\bf k}\over N^{2}}\sum_{\bf pq}\Omega^{({\rm s})}_{{\bf k}{\bf p}{\bf q}}{F^{\rm(s)}({\bf k},{\bf p},{\bf q})\over[\omega_{{\bf q}+{\bf k}}-(\bar{\xi}_{{\bf p}+{\bf q}}-\bar{\xi}_{{\bf p}})]^{4}},~{}~{}~{}~{}~{}~{}~{}~{}$ (16b)
respectively. Substituting the above spin self-energy (15) into Eq. (12), the full spin propagator in the normal-state can be derived as,
$D({\bf k},\omega)={\bar{B}_{1{\bf k}}\over\omega^{2}-\bar{\omega}^{2}_{1{\bf k}}}+{\bar{B}_{2{\bf k}}\over\omega^{2}-\bar{\omega}^{2}_{2{\bf k}}}=\sum_{\alpha=1,2}{\bar{B}_{\alpha{\bf k}}\over\omega^{2}-\bar{\omega}^{2}_{\alpha{\bf k}}},~{}~{}~{}~{}~{}~{}$ (17)
with the renormalized spin excitation energy dispersions,
$\bar{\omega}^{2}_{1{\bf k}}$ $=$ ${1\over 2}\left[\omega^{2}_{\bf k}+\omega^{2}_{0{\bf k}}+\sqrt{(\omega^{2}_{\bf k}-\omega^{2}_{0{\bf k}})^{2}+4B^{2}_{\bf k}[\bar{\Delta}^{({\rm s})}_{\rm pg}({\bf k})]^{2}}\right],$
$\bar{\omega}^{2}_{2{\bf k}}$ $=$ ${1\over 2}\left[\omega^{2}_{\bf k}+\omega^{2}_{0{\bf k}}-\sqrt{(\omega^{2}_{\bf k}-\omega^{2}_{0{\bf k}})^{2}+4B^{2}_{\bf k}[\bar{\Delta}^{({\rm s})}_{\rm pg}({\bf k})]^{2}}\right],$
and the corresponding weight functions of the full spin excitation spectrum,
$\bar{B}_{1{\bf k}}$ $=$ ${1\over 2}B_{\bf k}\left({\omega^{2}_{\bf k}-\omega^{2}_{0{\bf k}}\over\sqrt{(\omega^{2}_{\bf k}-\omega^{2}_{0{\bf k}})^{2}+4B^{2}_{\bf k}[\bar{\Delta}^{({\rm s})}_{\rm pg}({\bf k})}]^{4}}+1\right),~{}~{}~{}~{}$ (19a)
$\bar{B}_{2{\bf k}}$ $=$ $-{1\over 2}B_{\bf k}\left({\omega^{2}_{\bf k}-\omega^{2}_{0{\bf k}}\over\sqrt{(\omega^{2}_{\bf k}-\omega^{2}_{0{\bf k}})^{2}+4B^{2}_{\bf k}[\bar{\Delta}^{({\rm s})}_{\rm pg}({\bf k})}]^{4}}-1\right).~{}~{}~{}~{}~{}~{}~{}~{}$ (19b)
The momentum dependence of the spin pseudogap $\bar{\Delta}^{({\rm s})}_{\rm pg}({\bf k})$ in Eq. (15) evolves strongly with doping and temperature, as that in the charge-carrier pseudogap Feng15 ; Feng12 . On the other hand, weMa23 have shown that the umklapp scattering between electrons by the exchange of the MF effective spin propagator (7) can give a consistent description of the low-temperature T-linear resistivity in the overdoped strange-metal phase, where the dominant contribution to the scattering rate mainly comes from the antinodal umklapp scattering. In this case, we focus on the exotic features of the doping and temperature dependent spin pseudogap $\bar{\Delta}^{({\rm s})}_{\rm pg}({\bf k})$ at around the antinodal region for a convenience in the following discussions of the low-temperature electrical transport in the underdoped pseudogap phase.
Figure 2: (a) The spin pseudogap $\bar{\Delta}^{({\rm s})}_{\rm pg}({\bf k}_{\rm AN})$ at the antinode for temperature $T=0.002J$ and (b) the corresponding antinodal spin-pseudogap crossover temperature $T^{*}_{\rm s}$ as a function of doping, where ${\bf k}_{\rm AN}$ is the wave vector at the antinode.
In Fig. 2a, we plot $\bar{\Delta}^{({\rm s})}_{\rm pg}({\bf k}_{\rm AN})$ at the antinode as a function of doping with temperature $T=0.002J$, where ${\bf k}_{\rm AN}$ is the wave vector at the antinode of the Brillouin zone (BZ). Apparently, this antinodal spin pseudogap is insensitive to the doping in the slightly underdoped region, and then it falls off rapidly as the doping is grown in the heavily underdoped region. To our big surprise, this antinodal spin pseudogap abruptly disappears at around the optimal doping. This antinodal spin pseudogap formation in the underdoped regime lowers the density of the spin excitations in response to the intense electron umklapp scattering from the spin excitations associated with the antinodes. However, in the overdoped regime, the main properties of the antinodal spin excitations can be well described by the MF spin propagator (9). Furthermore, for a given doping, this antinodal spin pseudogap vanishes when temperature reaches the antinodal spin-pseudogap crossover temperature $T^{*}_{\rm s}$. To see this doping dependence of $T^{*}_{\rm s}$ more clearly, we plot $T^{*}_{\rm s}$ as a function of doping in Fig. 2b, where in corresponding to the result of $\bar{\Delta}^{({\rm s})}_{\rm pg}({\bf k}_{\rm AN})$ in Fig. 2a, $T^{*}_{\rm s}$ is relatively high in the slightly underdoped region, and then it decreases rapidly when the doping is increased in the heavily underdoped region, eventually disappearing at around the optimal doping.
Based on the full spin propagator (17) [then the dynamical spin structure factor], the dynamical spin response of cuprate superconductors in the normal-state has been investigatedKuang15 ; Yuan01 ; Feng98 , where the obtained results show the existence of damped but well-defined dispersive spin excitations in the whole doping phase diagram. Moreover, the low-energy spin fluctuation is dominated by the process from the mobile charge carriers, while the high-energy spin excitation on the other hand retains roughly constant energy as a function of doping, with spectral weight and dispersion relation comparable to those in the corresponding SC-state. All these results are qualitatively consistent with the experimental observations Fujita12 ; Birgeneau89 ; Fong95 ; Yamada98 ; Arai99 ; Bourges00 ; He01 ; Hayden04 ; Tranquada04 ; Bourges05 .
II.3 Electronic structure in the underdoped pseudogap state
For the understanding of the nature of the electronic structure in cuprate superconductors, it is needed to derive the electron propagator, which is characterized by the charge-spin recombination. In order to obtain the electron propagator, the full charge-spin recombination scheme has been developedFeng15a , where it has been shown that the coupling form between the electron and spin excitation in the $t$-$J$ model in the normal-state is the same as that between the charge carrier and spin excitation in Eq. (4), and then the self-consistent equation satisfied by the full electron propagator in the normal-state can be derived directly asFeng15a ,
$G({\bf k},\omega)=G^{(0)}({\bf k},\omega)+G^{(0)}({\bf k},\omega)\Sigma_{\rm ph}({\bf k},\omega)G({\bf k},\omega),~{}~{}~{}~{}$ (20)
where $G^{(0)}({\bf k},\omega)$ is the electron propagator of the $t$-$J$ model (1) in the tight-binding approximation, and has been derived as $G^{(0)-1}({\bf k},\omega)=\omega-\varepsilon_{\bf k}$ with the electron energy dispersion $\varepsilon_{\bf k}=-4t\gamma_{\bf k}+4t^{\prime}\gamma_{\bf k}^{\prime}+\mu$. In this case, the full electron propagator $G({\bf k},\omega)$ in Eq. (20) can be expressed explicitly as,
$G({\bf k},\omega)={1\over\omega-\varepsilon_{\bf k}-\Sigma_{\rm ph}({\bf k},\omega)},$ (21)
with the electron normal self-energy,
$\Sigma_{\rm ph}({\bf k},i\omega_{n})$ $=$ ${1\over N}\sum_{\bf p}{1\over\beta}\sum_{ip_{m}}G({\bf p}+{\bf k},ip_{m}+i\omega_{n})$ (22)
$\times$ $P^{(0)}({\bf k},{\bf p},ip_{m}),~{}~{}~{}$
which has been evaluated asFeng15a ,
$\Sigma_{\rm ph}({\bf k},\omega)$ $=$ ${1\over N^{2}}\sum_{{\bf pp^{\prime}}\mu\nu}(-1)^{\nu+1}Z_{\rm F}\Omega_{\bf pp^{\prime}k}$ (23)
$\times$ ${F_{\mu\nu}({\bf p},{\bf p}^{\prime}{\bf k})\over\omega+(-1)^{\mu+1}\omega_{\nu{\bf p}{\bf p}^{\prime}}-\bar{\varepsilon}_{{\bf p}+{\bf k}}},~{}~{}~{}$
with the function,
$F_{\mu\nu}({\bf p},{\bf p}^{\prime}{\bf k})=n_{\rm F}[(-1)^{\mu+1}\bar{\varepsilon}_{{\bf p}+{\bf k}}]n^{(\nu)}_{{\rm 1B}{\bf pp^{\prime}}}+n^{(\nu)}_{{\rm 2B}{\bf pp^{\prime}}},~{}~{}~{}~{}~{}$ (24)
where $\mu~{}(\nu)=1,~{}2$, $\bar{\varepsilon}_{\bf k}=Z_{\rm F}\varepsilon_{\bf k}$ is the renormalized electron energy dispersion, $Z^{-1}_{\rm F}=1-{\rm Re}\Sigma_{\rm pho}({\bf k},0)\mid_{{\bf k}=[\pi,0]}$ is the electron coherent weight, $\Sigma_{\rm pho}({\bf k},\omega)$ is the antisymmetric part of the electron normal self-energy $\Sigma_{\rm ph}({\bf k},\omega)$, $n^{(\nu)}_{{\rm 1B}{\bf pp^{\prime}}}=1+n_{\rm B}(\omega_{{\bf p}^{\prime}+{\bf p}})+n_{\rm B}[(-1)^{\nu+1}\omega_{\bf p^{\prime}}]$, and $n^{(\nu)}_{{\rm 2B}{\bf pp^{\prime}}}=n_{\rm B}(\omega_{{\bf p}^{\prime}+{\bf p}})n_{\rm B}[(-1)^{\nu+1}\omega_{\bf p^{\prime}}]$. This electron normal self-energy characterizes a competition between the kinetic energy and magnetic energy in the $t$-$J$ model (1).
The electron normal self-energy (23) originated from the charge-carrier normal self-energy (10) is due to the charge-spin recombination, indicating that the normal-state pseudogap state originated from the charge-carrier pseudogap state is due to the charge-spin recombination. To see the nature of the normal-state pseudogap more clearly, the above electron normal self-energy in Eq. (23) can be rewritten as,
$\Sigma_{\rm ph}({\bf k},\omega)$ $=$ ${[\bar{\Delta}_{\rm PG}({\bf k})]^{2}\over\omega-\varepsilon_{0{\bf k}}},$ (25)
where $\varepsilon_{0{\bf k}}=L_{2{\bf k}}/L_{1{\bf k}}$ is the energy spectrum of $\Sigma_{\rm ph}({\bf k},\omega)$, and $\bar{\Delta}^{2}_{\rm PG}({\bf k})=L^{2}_{2{\bf k}}/L_{1{\bf k}}$ is the normal-state pseudogap, with the functions $L_{1}({\bf k})=-\Sigma_{\rm pho}({\bf k},\omega=0)$ and $L_{2}({\bf k})=-\Sigma_{\rm ph}({\bf k},\omega=0)$, while $\Sigma_{\rm pho}({\bf k},\omega=0)$ and the antisymmetric part $\Sigma_{\rm pho}({\bf k},\omega)$ of the electron normal self-energy can be derived directly from the electron normal self-energy $\Sigma_{\rm ph}({\bf k},\omega)$ in Eq. (23). This normal-state pseudogap $\bar{\Delta}_{\rm PG}({\bf k})$ is also identified as being a region of the electron normal self-energy in which the normal-state pseudogap anisotropically suppresses the electronic density of states on EFS. Since the normal-state pseudogap state is induced by the charge-carrier pseudogap state, the normal-state pseudogap $\bar{\Delta}_{\rm PG}({\bf k})$ [the normal-state pseudogap crossover temperature $T^{*}$] as a function of doping presents a similar behavior of the charge-carrier pseudogap $\bar{\Delta}^{({\rm h})}_{\rm pg}({\bf k})$ [the charge-carrier pseudogap crossover temperature $T^{*}_{\rm h}$]Feng15 ; Feng12 . This normal-state pseudogap crossover temperature $T^{*}$ is actually a crossover line below which a novel electronic state emerges, as exemplified by the presence of the Fermi arcs, the competing electronic orders, etc., and then the unconventional features of this novel electronic state can be well interpreted in terms of the formation of the normal-state pseudogap Damascelli03 ; Campuzano04 ; Fink07 ; Deutscher05 ; Devereaux07 ; Fischer07 ; Hussey08 ; Kastner98 ; Timusk99 ; Hufner08 ; Vishik18 .
The measured energy and momentum distribution curves can be depicted by the electron excitation spectrum Dessau91 ; Randeria95 ; Fedorov99 ; Campuzano99 ; DMou17 ; Bogdanov00 ; Kaminski01 ; Johnson01 ; Iwasawa08 ; Plumb13 ,
$I({\bf k},\omega)\propto n_{\rm F}(\omega)A({\bf k},\omega),~{}~{}~{}~{}$ (26)
with the electron spectral function,
$A({\bf k},\omega)={1\over\pi}{{\rm Im}\Sigma_{\rm ph}({\bf k},\omega)\over[\omega-\varepsilon_{\bf k}-{\rm Re}\Sigma_{\rm ph}({\bf k},\omega)]^{2}+[{\rm Im}\Sigma_{\rm ph}({\bf k},\omega)]^{2}},~{}~{}~{}$ (27)
where ${\rm Re}\Sigma_{\rm ph}({\bf k},\omega)$ and ${\rm Im}\Sigma_{\rm ph}({\bf k},\omega)$ are the real and imaginary parts of the electron normal self-energy $\Sigma_{\rm ph}({\bf k},\omega)$, respectively. The electrons are renormalzed due to the electron scattering mediated by the spin excitation and then they acquire a finite lifetime.
The EFS topology is known to be crucial in the understanding of the unconventional superconductivity in cuprate superconductors Damascelli03 ; Campuzano04 ; Fink07 ; Deutscher05 ; Devereaux07 ; Fischer07 and the related anomalous normal-state properties Hussey08 ; Kastner98 ; Timusk99 ; Hufner08 ; Vishik18 . The ARPES measured EFS can be described theoretically by the intensity map of the electron excitation spectrum (26) at zero energy $\omega=0$, where the locations of the EFS continuous contour in momentum space is determined directly by the poles of the electron spectral function (27),
Figure 3: (Color online) Electron Fermi surface at $\delta=0.09$ with $T=0.002J$, where the Brillouin zone center has been shifted by [$\pi,\pi$], and AN, TFA, and ND denote the antinode, tip of the Fermi arc, and node, respectively.
$\varepsilon_{\bf k}+{\rm Re}\Sigma_{\rm ph}({\bf k},0)=\bar{\varepsilon}_{\bf k}=0$, while the strong redistribution of the spectral weight on EFS is mainly governed by the momentum dependence of the normal-state pseudogap $\bar{\Delta}_{\rm PG}({\bf k})$ [then the single-electron scattering rate $\Gamma_{\bf k}(\omega)={\rm Im}\Sigma_{\rm ph}({\bf k},\omega)=\pi[\bar{\Delta}_{\rm PG}({\bf k})]^{2}\delta(\omega+\varepsilon_{0{\bf k}})]$. For a convenience in the following discussions of the electrical transport in the underdoped pseudogap phase, the EFS mapFeng16 ; Liu21 ; Cao21 ; Zeng22 at doping $\delta=0.09$ with temperature $T=0.002J$ is replotted in Fig. 3, where the BZ center has been shifted by [$\pi,\pi$], and AN, TFA, and ND indicate the antinode, tip of the Fermi arc, and node, respectively. It thus shows clearly that the antinodal region of EFS becomes partially gapped, leading to that EFS consists, not of a closed contour, but only of four disconnected Fermi arcs centered around the nodes, in qualitative agreement with the experimental results Loeser96 ; Norman98 ; Chatterjee06 ; Shi08 ; Sassa11 ; Kaminski15 ; Loret17 ; Loret18 ; He14 ; Comin14 ; Comin16 . However, the most of the spectral weight on the Fermi arcs locates at around the tips of the Fermi arcs, indicating that the electrons at around the tips of the Fermi arcs have a largest density of states Chatterjee06 ; Shi08 ; Sassa11 ; Kaminski15 ; Loret17 ; Loret18 ; He14 ; Comin14 ; Comin16 , and then charge order is driven by this EFS instability, with a characteristic wave vector corresponding to the tips of the Fermi arcsFeng16 ; Comin14 ; Comin16 .
The elementary excitations are parameterized by the electron spectral function (27), which has given a consistent description of the renormalization of the electrons in the underdoped pseudogap phaseFeng16 ; Liu21 ; Cao21 ; Zeng22 , where the obtained results show that the single-electron scattering rate has a well-pronounced peak structure at around the antinodal and nodal regions, which leads to the remarkable peak-dip-hump structure in the energy distribution curve. Moreover, the striking dispersion kink is induced by the inflection point in the single-electron scattering rate, while the spectral weight at around the dispersion kink is reduced highly by the corresponding peak in the real part of the electron normal self-energy. All these results are in qualitative agreement with the corresponding experimental results Dessau91 ; Randeria95 ; Fedorov99 ; Campuzano99 ; DMou17 ; Bogdanov00 ; Kaminski01 ; Johnson01 ; Iwasawa08 ; Plumb13 . More importantly, the anisotropic suppression of the electronic density of states on EFS by the normal-state pseudogap can affect the electrical transport in two ways Timusk99 : (i) through the reduction of the number of electron current-carrying states as a normal-state pseudogap forms; and (ii) since the electron current carriers are scattered by the spin excitation, through the reduction in the density of spin excitations.
II.4 Electrical transport due to umklapp scattering from a spin excitation
We now turn to discuss the low-temperature electrical transport in the underdoped pseudogap phase. Theoretically, the Boltzmann transport equation is the cornerstone for the discussions of the electrical transportAbrikosov88 ; Mahan81 , since the Boltzmann transport equation is valid in the case of either the existence of the well-defined quasiparticles or the treatment of the electron interaction mediated by different bosonic modes within the Eliashberg approach. This builds on the following pioneering works: (i) in the early days of the electrical transport research in the conventional superconductorsPrange64 , a set of the transport equations has been derived in the Migdal’s approximation, where this coupled set of the transport equations for the electron and phonon distribution functions are correct even in the case of the absence of the well-defined quasiparticles. In particular, one of the forms of the electrical transport equation,
$e{\bf E}\cdot\nabla_{\bf k}f({\bf k})=I_{\rm e-e},$ (28)
is identical to the electrical Boltzmann transport equation suggested by Landau for the case in which the quasiparticle is well-definedAbrikosov88 ; Mahan81 , with the charge of an electron $e$, the electron distribution function in a homogeneous system $f({\bf k},t)$, and the electron-electron collision term $I_{\rm e-e}$. For a convenience in the following discussions, the external magnetic field ${\bf H}$ has been dropped, and only an external electric field ${\bf E}$ is applied to the system; (ii) however, this Boltzmann transport equation (28) is not specific to the electron interaction mediated by the phonon in the conventional superconductors, and has been confirmed recently that it is also valid for the system with the electron interaction mediated by other bosonic excitations Lee21 ; Ma23 . With the advantage of their insightLee21 ; Ma23 ; Prange64 , the electrical transport due to the electron scattering mediated by various kinds of bosonic modes can be evaluated in a simple way even in the case of the break down of the quasiparticle picture.
In the following discussions, we study the the low-temperature electrical transport in the underdoped pseudogap phase based on the electrical Boltzmann transport equation (28). For the calculation of this electrical Boltzmann transport equation (28), the linear perturbation from the equilibrium in terms of the distribution function has been introduced Lee21 ; Ma23 ; Prange64 ,
$f({\bf k})$ $=$ $n_{\rm F}({\bar{\varepsilon}_{\bf k}})-{dn_{\rm F}({\bar{\varepsilon}_{\bf k}})\over d{\bar{\varepsilon}_{\bf k}}}\tilde{\Phi}({\bf k}),$ (29)
with the local shift of the chemical potential at a given patch of EFS $\tilde{\Phi}({\bf k})$, which satisfied the antisymmetric relation $\tilde{\Phi}(-{\bf k})=-\tilde{\Phi}({\bf k})$.
Figure 4: (Color online) Schematic picture of the electron umklapp scattering process, where an electron on a electron Fermi surface (left) is scattered by its partner on the umklapp electron Fermi surface (right). The Fermi wave vector of the tips of the Fermi arcs ${\rm k}^{\rm TFA}_{\rm F}$ is the radius of the circular electron Fermi surface (red), and then an electron on this circular electron Fermi surface (left) parametrized by the Fermi angle $\theta$ is scattered to a point parametrized by the Fermi angle $\theta^{\prime}$ on the umklapp electron Fermi surface (right) by the spin excitation carrying momentum ${\rm p}(\theta,\theta^{\prime})$. $\Delta_{p}$ is the minimal umklapp vector at the antinode (the Fermi angle $\theta=0$).
We substitute the above result in Eq. (29) into Eq. (28), and then linearize the electrical Boltzmann equation (28) as,
$e{\bf v}_{\bf k}\cdot{\bf E}{dn_{\rm F}({\bar{\varepsilon}_{\bf k}})\over d{\bar{\varepsilon}_{\bf k}}}=I_{\rm e-e},$ (30)
where ${\bf v}_{\bf k}=\nabla_{\bf k}{\bar{\varepsilon}_{\bf k}}$ is the electron velocity.
The electron-electron collision term $I_{\rm e-e}$ in the electrical Boltzmann equation is directly connected with the electrical scattering mechanism Abrikosov88 ; Mahan81 . Although there is no consensus on the origin of the low-temperature electrical transport in the underdoped pseudogap phase to date, it is widely believed that the opening of the pseudogap induces a deviation from the low-temperature T-linear behaviour of the resistivity. In particular, it has been argued that if the electron scattering responsible for the low-temperature T-linear resistivity in the overdoped regime involves the electron scattering on the spin excitations in the underdoped regimePuchkov96 , then the opening of the spin pseudogap may naturally lead to a deviation from the low-temperature T-linear behaviour of the resistivityBucher93 ; Ito93 ; Nakano94 ; Ando01 . On the other hand, it has been shown clearly that the electron umklapp scattering is the origin of the low-temperature T-linear resistivity in the strange-metal phase Lee21 ; Honerkamp01 ; Hartnoll12 ; Ma23 . Moreover, weMa23 have shown very recently that the low-temperature T-linear resistivity in the overdoped strange-metal phase arises from the umklapp scattering between electrons by the exchange of the MF effective spin propagator, where the low-temperature electronic transport is mainly governed by the antinodal umklapp scattering, and then the low-temperature umklapp scattering rate is T-linear near the umklapp point, which naturally generates a low-temperature T-linear resistivity. In this paper, we employ the electron umklapp scattering to study the low-temperature electrical transport in the underdoped pseudogap phase by considering the spin pseudogap effects. To see the electron umklapp scattering process more clearly, we show a schematic picture of the electron umklapp scattering processLee21 ; Ma23 in Fig. 4, where an electron on a circular EFS (left) is scattered by its partner on the umklapp EFS (right). It should be noted that the intensity map of EFS in Fig. 4 is the same as the EFS map in Fig. 3, while the Fermi wave vector of the tips of the Fermi arcs ${\rm k}^{\rm TFA}_{\rm F}$ is the radius of the EFS circle (red). In particular, this circle EFS (red) connects all tips of the Fermi arcs, and then the most of the electronic density of states is concentred on this circular EFS.
Following these recent discussionsLee21 ; Ma23 , we can derive the electron-electron collision $I_{\rm e-e}$ in Eq. (30) in the underdoped pseudogap phase as,
$I_{\rm e-e}$ $=$ ${1\over N^{2}}\sum_{{\bf k}^{\prime},{\bf p}}{2\over T}|P({\bf k},{\bf p},{\bf k}^{\prime},\bar{\varepsilon}_{\bf k}-\bar{\varepsilon}_{{\bf k}+{\bf p}+{\bf G}})|^{2}\{\tilde{\Phi}({\bf k})+\tilde{\Phi}({\bf k^{\prime}})-\tilde{\Phi}({\bf k}+{\bf p}+{\bf G})-\tilde{\Phi}({\bf k}^{\prime}-{\bf p})\}$ (31)
$\times$ $n_{\rm F}(\bar{\varepsilon}_{\bf k})n_{\rm F}(\bar{\varepsilon}_{{\bf k}^{\prime}})[1-n_{\rm F}(\bar{\varepsilon}_{{\bf k}+{\bf p}+{\bf G}})][1-n_{\rm F}(\bar{\varepsilon}_{{\bf k}^{\prime}-{\bf p}})]\delta(\bar{\varepsilon}_{\bf k}+\bar{\varepsilon}_{\bf k^{\prime}}-\bar{\varepsilon}_{{\bf k}+{\bf p}+{\bf G}}-\bar{\varepsilon}_{{\bf k}^{\prime}-{\bf p}}),$
where ${\bf G}$ represents a set of reciprocal lattice vectors. It should be emphasized that the above electron umklapp scattering (31) is described as a scattering between electrons by the exchange of the effective spin propagator,
$P({\bf k},{\bf p},{\bf k}^{\prime},\omega)$ $=$ ${1\over N}\sum_{\bf q}\Lambda_{{\bf p}+{\bf q}+{\bf k}}\Lambda_{{\bf q}+{\bf k}^{\prime}}\bar{\Pi}({\bf p},{\bf q},\omega),~{}~{}~{}$ (32)
rather than the scattering between electrons via the emission and absorption of the spin excitationLee21 ; Ma23 .
In cuprate superconductors, a small density of charge carriers is sufficient to destroy the AF long-range order. However, the doped charge carriers and magnetic-exchange-coupled spins organize themselves in a cooperative way to enhance both the electrons mobility and the AF short-range order correlation, and then the spin excitations in the spin liquid state with the AF short-range order appear to survive from the underdoped regime to the overdoped regimeFujita12 . Moreover, it has been shown that the coupling strength of the electrons with the spin excitations gradually weakens with the increase of doping from a strong-coupling case in the underdoped regime to a weak-coupling side in the overdoped regime Feng12 ; Johnson01 ; Kordyuk10 , reflecting a reduction of the strength of the magnetic fluctuation with the increase of doping. In other words, (i) the effect of the magnetic fluctuation in the overdoped regime is less dramatic than in the underdoped regimeKastner98 ; Timusk99 ; Hufner08 ; Vishik18 ; Fujita12 ; (ii) on the other hand, as shown in Fig. 2, the effect from antinodal spin pseudogap is absent in the overdoped regime, and then the main properties of the antinodal spin excitations can be well described by the MF spin propagator (9). These are reasons why the interpretation of the low-temperature T-linear resistivity in the overdoped strange-metal phase can be well made in terms of the umklapp scattering between electrons by the exchange of the MF effective spin propagatorMa23 .
However, in the strong-coupling side (then in the underdoped regime), (i) the effect of the magnetic fluctuation is much dramatic Kastner98 ; Timusk99 ; Hufner08 ; Vishik18 ; Fujita12 ; and (ii) as shown in Fig. 2, the effect from antinodal spin pseudogap is particularly notable, where the antinodal spin pseudogap lowers the density of the spin excitations in response to the intense electron umklapp scattering, which reduces the strength of the electron umklapp scattering from the spin excitation states into the antinodal region. In this case, the umklapp scattering between electrons should be mediated by the exchange of the full effective spin propagator for a proper description of the low-temperature electrical transport in the underdoped pseudogap phase.
Now our aim is to obtain the full effective spin propagator. The full spin bubble $\bar{\Pi}({\bf p},{\bf q},\omega)$ in Eq. (32) is derived directly in terms of the full spin propagator (17) as,
$\bar{\Pi}({\bf p},{\bf q},\omega)=-\sum_{\begin{subarray}{c}\alpha=1,2\\ \alpha^{\prime}=1,2\end{subarray}}{\bar{W}^{(1)}_{\alpha\alpha^{\prime}{\bf p}{\bf q}}\over\omega^{2}-[\bar{\omega}^{(1)}_{\alpha\alpha^{\prime}{\bf p}{\bf q}}]^{2}}+{\bar{W}^{(2)}_{\alpha\alpha^{\prime}{\bf p}{\bf q}}\over\omega^{2}-[\bar{\omega}^{(2)}_{\alpha\alpha^{\prime}{\bf p}{\bf q}}]^{2}},~{}~{}~{}~{}~{}~{}$ (33)
where the spin excitation energy dispersions $\bar{\omega}^{(1)}_{\alpha\alpha^{\prime}{\bf p}{\bf q}}$ and $\bar{\omega}^{(2)}_{\alpha\alpha^{\prime}{\bf p}{\bf q}}$ are given by,
$\bar{\omega}^{(1)}_{\alpha\alpha^{\prime}{\bf p}{\bf q}}$ $=$ $\bar{\omega}_{\alpha{\bf q}+{\bf p}}+\bar{\omega}_{\alpha^{\prime}{\bf q}},~{}~{}~{}~{}$ (34a)
$\bar{\omega}^{(2)}_{\alpha\alpha^{\prime}{\bf p}{\bf q}}$ $=$ $\bar{\omega}_{\alpha{\bf q}+{\bf p}}-\bar{\omega}_{\alpha^{\prime}{\bf q}},$ (34b)
respectively, and the corresponding functions,
$\bar{W}^{(1)}_{\alpha\alpha^{\prime}{\bf p}{\bf q}}$ $=$ ${\bar{B}_{\alpha^{\prime}{\bf q}}\bar{B}_{\alpha{\bf q}+{\bf p}}\over 2\bar{\omega}_{\alpha^{\prime}{\bf q}}\bar{\omega}_{\alpha{\bf q}+{\bf p}}}\bar{\omega}^{(1)}_{\alpha\alpha^{\prime}{\bf p}{\bf q}}$ (35)
$\times$ $[n_{\rm B}(\bar{\omega}_{\alpha{\bf q}+{\bf p}})+n_{\rm B}(\bar{\omega}_{\alpha^{\prime}{\bf q}})+1],~{}~{}~{}~{}~{}$
$\bar{W}^{(2)}_{\alpha\alpha^{\prime}{\bf p}{\bf q}}$ $=$ ${\bar{B}_{\alpha^{\prime}{\bf q}}\bar{B}_{\alpha{\bf q}+{\bf p}}\over 2\bar{\omega}_{\alpha^{\prime}{\bf q}}\bar{\omega}_{\alpha{\bf q}+{\bf p}}}\bar{\omega}^{(2)}_{\alpha\alpha^{\prime}{\bf p}{\bf q}}$ (36)
$\times$ $[n_{\rm B}(\bar{\omega}_{\alpha{\bf q}+{\bf p}})-n_{\rm B}(\bar{\omega}_{\alpha^{\prime}{\bf q}})].$
The full effective spin propagator $P({\bf k},{\bf p},{\bf k}^{\prime},\omega)$ in Eq. (32) now can be derived from the above spin bubble (33) as,
$P({\bf k},{\bf p},{\bf k}^{\prime},\omega)$ $=$ $-{1\over N}\sum\limits_{\alpha\alpha^{\prime}{\bf q}}\left[{\varpi^{(1)}_{\alpha\alpha^{\prime}}({\bf k},{\bf p},{\bf k}^{\prime},{\bf q})\over\omega^{2}-[\bar{\omega}^{(1)}_{\alpha\alpha^{\prime}{\bf p}{\bf q}}]^{2}}\right.$ (37)
-- $\left.{\varpi^{(2)}_{\alpha\alpha^{\prime}}({\bf k},{\bf p},{\bf k}^{\prime},{\bf q})\over\omega^{2}-[\bar{\omega}^{(2)}_{\alpha\alpha^{\prime}{\bf p}{\bf q}}]^{2}}\right],~{}~{}~{}~{}~{}$
with the weight functions,
$\varpi^{(1)}_{\alpha\alpha^{\prime}}({\bf k},{\bf p},{\bf k}^{\prime},{\bf q})$ $=$ $\Lambda_{{\bf k}+{\bf p}+{\bf q}}\Lambda_{{\bf q}+{\bf k}^{\prime}}\bar{W}^{(1)}_{\alpha\alpha^{\prime}{\bf p}{\bf q}},$ (38a)
$\varpi^{(2)}_{\alpha\alpha^{\prime}}({\bf k},{\bf p},{\bf k}^{\prime},{\bf q})$ $=$ $\Lambda_{{\bf k}+{\bf p}+{\bf q}}\Lambda_{{\bf q}+{\bf k}^{\prime}}\bar{W}^{(2)}_{\alpha\alpha^{\prime}{\bf p}{\bf q}}.$ (38b)
The electron umklapp scattering in Eq. (31 shows that the electron-electron collision $I_{\rm e-e}$ is both functions of momentum and energy. Howerver, the amplitude of the wave vector dependence is unimportant at low temperatures, since everything happens at EFSAbrikosov88 ; Mahan81 . In this case, an any given patch on the circular EFS shown in Fig. 4 is represented via the Fermi angle $\theta$ with the Fermi angle range $\theta\in[0,2\pi]$, and then the momentum integration along the perpendicular momentum can be replaced by the integrationLee21 ; Ma23 ; Prange64 over $\bar{\varepsilon}_{\bf k}$. For the umklapp scattering between electrons by the exchange of the full effective spin propagator in Eq. (31), an electron on the circular EFS parametrized by the Fermi angle $\theta$ is scattered to a point parametrized by the Fermi angle $\theta^{\prime}$ on the umklapp EFS via the spin excitation carrying momentum ${\rm p}(\theta,\theta^{\prime})$ as shown in Fig. 4. Following the above treatment and the calculation process in the recent worksLee21 ; Ma23 for the low-temperature electrical transport in the overdoped strange-metal phase, the electron-electron collision $I_{\rm e-e}$ in Eq. (31) in the underdoped pseudogap phase can be obtained straightforwardlyMa23 , and then the electrical Boltzmann transport equation (30) can be derived as,
$e{\bf v}_{\rm F}(\theta)\cdot{\bf E}=-2\int{d\theta^{\prime}\over{2\pi}}\zeta(\theta^{\prime})F(\theta,\theta^{\prime})[\Phi(\theta)-\Phi(\theta^{\prime})],~{}~{}~{}~{}~{}$ (39)
with $\Phi(\theta)=\tilde{\Phi}[{\rm k}(\theta)]$, the Fermi velocity ${\bf v}_{\rm F}(\theta)$ at the Fermi angle $\theta$, the density of states factor $\zeta(\theta^{\prime})={\rm k}^{2}_{\rm F}/[4\pi^{2}{\rm v}^{3}_{\rm F}]$ at angle $\theta^{\prime}$, the Fermi wave vector ${\rm k}_{\rm F}$, and the Fermi velocity ${\rm v}_{\rm F}$. In particular, the antisymmetric relation $\tilde{\Phi}(-{\bf k})=-\tilde{\Phi}({\bf k})$ for $\tilde{\Phi}({\bf k})$ in Eq. (29) has been converted into $\Phi(\theta)=-\Phi(\theta+\pi)$ for $\Phi(\theta)$ in the above Eq. (39). Moreover, the coefficient of $\Phi(\theta)$ in the first term of the right-hand side of Eq. (39),
$\gamma(\theta)=2\int{d\theta^{\prime}\over{2\pi}}\zeta(\theta^{\prime})F(\theta,\theta^{\prime}),$ (40)
can be referred to as the angular (momentum) dependence of the umklapp scattering rateMa23 ; Lee21 , with the kernel function $F(\theta,\theta^{\prime})$ that connects the point $\theta$ on the circular EFS with the point $\theta^{\prime}$ on the umklapp EFS via the amplitude of the momentum transfer ${\rm p}(\theta,\theta^{\prime})$ as shown in Fig. 4, which can be derived as,
$F(\theta,\theta^{\prime})$ $=$ ${1\over T}\int{d\omega\over 2\pi}{\omega^{2}\over{\rm p}(\theta,\theta^{\prime})}{|\bar{P}[{\rm k}(\theta),{\rm p}(\theta,\theta^{\prime}),\omega]|}^{2}$ (41)
$\times$ $n_{\rm B}(\omega)[1+n_{\rm B}(\omega)],~{}~{}~{}~{}~{}~{}$
where the full effective spin propagator $\bar{P}[{\rm k}(\theta),{\rm p}(\theta,\theta^{\prime}),\omega]$ has been rewritten explicitly in terms of the Fermi angles $\theta$ and $\theta^{\prime}$ asMa23 ,
$\bar{P}[{\rm k}(\theta),{\rm p}(\theta,\theta^{\prime},\omega]$ $=$ $-{1\over N}\sum\limits_{\alpha\alpha^{\prime}{\bf q}}\left[{\varpi^{(1)}_{\alpha\alpha^{\prime}}(\theta,\theta^{\prime},{\bf q})\over\omega^{2}-[\bar{\omega}^{(1)}_{\alpha\alpha^{\prime}\theta,\theta^{\prime}}({\bf q})]^{2}}\right.$ (42)
-- $\left.{\varpi^{(2)}_{\alpha\alpha^{\prime}}(\theta,\theta^{\prime},{\bf q})\over\omega^{2}-[\bar{\omega}^{(2)}_{\alpha\alpha^{\prime}\theta,\theta^{\prime}}({\bf q})]^{2}}\right],~{}~{}~{}~{}~{}$
where $\varpi^{(1)}_{\alpha\alpha^{\prime}}(\theta,\theta^{\prime},{\bf q})=\varpi^{(1)}_{\alpha\alpha^{\prime}}[{\rm k}(\theta),{\rm p}(\theta,\theta^{\prime}),{\bf k}^{\prime}_{\rm F},{\bf q}]$, $\varpi^{(2)}_{\alpha\alpha^{\prime}}(\theta,\theta^{\prime},{\bf q})=\varpi^{(2)}_{\alpha\alpha^{\prime}}[{\rm k}(\theta),{\rm p}(\theta,\theta^{\prime}),{\bf k}^{\prime}_{\rm F},{\bf q}]$, $\bar{\omega}^{(1)}_{\alpha\alpha^{\prime}\theta,\theta^{\prime}}({\bf q})=\bar{\omega}^{(1)}_{\alpha\alpha^{\prime}{\rm p}(\theta,\theta^{\prime}){\bf q}}$, and $\bar{\omega}^{(2)}_{\alpha\alpha^{\prime}\theta,\theta^{\prime}}({\bf q})=\bar{\omega}^{(2)}_{\alpha\alpha^{\prime}{\rm p}(\theta,\theta^{\prime}){\bf q}}$.
For the discussions of the low-temperature resistivity, we need to obtain the electron current density, which can be derived in terms of the local shift of the chemical potential $\Phi(\theta)$ asMa23 ,
${\bf J}$ $=$ $en_{0}{1\over N}\sum_{\bf k}{\bf v}_{\bf k}{dn_{\rm F}({\bar{\varepsilon}_{\bf k}})\over d\bar{\varepsilon}_{\bf k}}\tilde{\Phi}({\bf k})$ (43)
$=$ $-en_{0}{{\rm k}_{\rm F}\over{\rm v}_{\rm F}}\int{d\theta\over(2\pi)^{2}}{\bf v}_{\rm F}(\theta)\Phi(\theta),~{}~{}~{}~{}~{}$
with the momentum relaxation that is generated by the action of the electric field on the mobile electrons at EFS with the density $n_{0}$. However, this local shift of the chemical potential $\Phi(\theta)$ can be evaluated directly in the relaxation-time approximation asLee21 ; Ma23 , $\Phi(\theta)=-e{\rm v}_{\rm F}{\rm cos}(\theta)E_{\hat{x}}/[2\gamma(\theta)]$, where the electric field ${\bf E}$ has been selected along the $\hat{x}$-axis. With the help of the above treatment, the dc conductivity can be derived as Lee21 ; Ma23 ,
$\sigma_{\rm dc}(T)={1\over 2}e^{2}n_{0}{\rm k}_{\rm F}{\rm v}_{\rm F}\int{d\theta\over(2\pi)^{2}}{\rm cos}^{2}(\theta){1\over\gamma(\theta)},$ (44)
and then the electrical resistivity can be expressed directly in terms of the above dc conductivity as,
$\rho(T)={1\over\sigma_{\rm dc}(T)}.$ (45)
It thus shows that the electrical resistance originated from the umklapp scattering between electrons by the exchange of the full effective spin propagator.
III Quantitative characteristics
Now we are ready to discuss the low-temperature electrical transport in the underdoped pseudogap phase. In Fig. 5, we plot the resistivity $\rho(T)$ in Eq. (45) as a function of temperature at doping $\delta=0.06$ (black-line), $\delta=0.09$ (red-line), and $\delta=0.12$ (blue-line), where in the low-temperature region, a power-law resistivity appears over a wide doping range in the underdoped regime.
Figure 5: (Color online) Resistivity as a function of temperature at $\delta=0.06$ (black-line), $\delta=0.09$ (red-line), and $\delta=0.12$ (blue-line). Inset: the numerical fit of the resistivity (black-dots) at $\delta=0.09$ with the fit form $\rho(T)=A_{2}T^{2}$.
Figure 6: (a) Strength of the T-quadratic resistivity as a function of doping. (b) The corresponding experimental result of cuprate superconductors taken from Ref. Barisic13, .
To explore this power-law behaviour of the low-temperature resistivity more clearly, the above results of $\rho(T)$ in Fig. 5 have been numerically fitted with the fit form $\rho(T)=A_{2}T^{2}$, and the fitted result of the low-temperature resistivity (black-dots) at $\delta=0.09$ is also plotted in Fig. 5 (inset), where the low-temperature resistivity is demonstrated clearly to be grown quadratically as the temperature is raised. Moreover, the strength of the T-quadratic resistivity (then the T-quadratic resistivity coefficient) $A_{2}$ grows as the doping is reduced. To see this doping dependence of the T-quadratic resistivity strength more clearly, we plot $A_{2}$ as a function of doping $\delta$ in Fig. 6a. For a better comparison, the corresponding transport experimental resultBarisic13 observed on the underdoped pseudogap phase of cuprate superconductors is also shown in Fig. 6b. It thus shows clearly that $A_{2}$ drops gradually as doping is enhanced in the underdoped regime, indicating that $A_{2}$ is roughly proportional to the inverse doping concentration. All these results are in qualitative agreement with the corresponding transport experimental results observed on the underdoped pseudogap phase of cuprate superconductors Ando04a ; Ando04 ; Lee05 ; Proust08 ; Cooper09 ; Hussey11 ; Barisic13 ; Pelc20 . These findings in Fig. 5 and Fig. 6 in the underdoped pseudogap phase along with the recent studyMa23 on the low-temperature T-linear resistivity in the overdoped strange-metal phase therefore show that the electron umklapp scattering from a spin excitation responsible for the T-linear resistivity in the overdoped regime naturally produces the low-temperature T-quadratic resistivity in the underdoped regime due to the opening of the momentum dependence of the spin pseudogap. As a nature consequence, the low-temperature T-linear resistivity in the overdoped strange-metal phase is dramatically switched to the low-temperature T-quadratic resistivity in the underdoped pseudogap phase.
Figure 7: (Color online) Surface plot of the kernel function $F(\theta,\theta^{\prime})$ at $\delta=0.09$ with $T=0.05J$, where AN, TFA, and ND denote the antinode, tip of the Fermi arc, and node, respectively.
The low-temperature resistivity is mainly determined by the umklapp scattering rate $\gamma(\theta,T)$ in Eq. (40), however, this umklapp scattering rate $\gamma(\theta,T)$ is in turn directly connected with the kernel function $F(\theta,\theta^{\prime})$ (then the probability weight of the umklapp scattering) in Eq. (41). For the further understanding of the nature of the umklapp scattering between electrons by the exchange of the full effective spin propagator, we first analyse the exotic feature of the kernel function $F(\theta,\theta^{\prime})$. In Fig. 7, we plot $F(\theta,\theta^{\prime})$ at $\delta=0.09$ with $T=0.05J$, where the probability weight of the electron umklapp scattering is very strongly anisotropic in momentum space. In particular, as in the case of the overdoped strange-metal phase Ma23 , the almost all of the probability weight of the electron umklapp scattering in the underdoped pseudogap phase is also concentrated at around the antinodal region, leading to that the strongest umklapp scattering occurs at around this antinodal region.
Figure 8: Umklapp scattering rate $\gamma(T)$ at the antinode as a function of temperature for $\delta=0.09$.
However, the weakest umklapp scattering appears at around the tips of the Fermi arcs, since the probability weight of the electron umklapp scattering there is almost none. Moreover, the strength of the electron umklapp scattering at around the nodal region is much weaker than that at around the antinodal region, since a very small amount of the probability weight of the electron umklapp scattering is concentrated at around the nodal region. This special momentum dependent distribution of the probability weight of the electron umklapp scattering in Fig. 7 together with the recent discussions of the momentum-dependent distribution of the probability weight of the electron umklapp scattering in the overdoped strange-metal phaseMa23 therefore confirm that in the whole doping regime, the dominant contribution to the scattering rate mainly arises from the umklapp scattering between the relatively slow electrons at around the antinodal region.
We now turn to explore the unusual evolution of the umklapp scattering rate (40) with temperature in the underdoped pseudogap phase. We have made a series of calculations for the umklapp scattering rate $\gamma(T)$ at different Fermi angles, and the result of $\gamma(T)$ as a function of temperature for $\delta=0.09$ at the antinode is plotted in Fig. 8, where the numerical fit demonstrates that $\gamma(T)$ is purely T-quadratic in the low-temperature region, i.e., it increases quadratically with temperature as the temperature is increased. More specifically, although the magnitude of $\gamma(\theta,T)$ is strongly anisotropic in momentum space, the low-temperature T-quadratic behaviour of $\gamma(\theta,T)$ appears at an any given Fermi angle $\theta$. Comparing this result in Fig. 8 with the corresponding results of the low-temperature T-quadratic resistivity in Fig. 5, it thus shows that the low-temperature T-quadratic behaviour of $\gamma(T)$ together with the temperature region are the same as the corresponding behaviour and region in the resistivity $\rho(T)$, which therefore confirms that the low-temperature T-quadratic resistivity with the corresponding temperature region is mainly governed by the low-temperature T-quadratic umklapp scattering rate with the corresponding temperature region.
Now we give an explanation to show why the low-temperature resistivity exhibits a dramatic switch from the T-linear resistivity in the overdoped strange-metal phase to the T-quadratic resistivity in the underdoped pseudogap phase? As the mentioned above, the nature of the umklapp scattering rate in Eq. (40) is mainly determined by the nature of the kernel function $F(\theta,\theta^{\prime})$ in Eq. (41). However, this kernel function $F(\theta,\theta^{\prime})$ in the underdoped pseudogap phase is proportional to the full effective spin propagator $P({\bf k},{\bf p},{\bf k}^{\prime},\omega)$ in Eq. (32). From the full spin propagator in Eq. (17), the full spin spectral function is derived straightforwardly as $A_{\rm spin}({\bf k},\omega)=A^{(1)}_{\rm spin}({\bf k},\omega)+A^{(2)}_{\rm spin}({\bf k},\omega)$, where the components $A^{(1)}_{\rm spin}({\bf k},\omega)$ and $A^{(2)}_{\rm spin}({\bf k},\omega)$ are given by,
$A^{(1)}_{\rm spin}({\bf k},\omega)$ $=$ ${\bar{B}_{1{\bf k}}\over\pi\bar{\omega}_{1{\bf k}}}[\delta(\omega-\bar{\omega}_{1{\bf k}})-\delta(\omega+\bar{\omega}_{1{\bf k}})],~{}~{}~{}~{}$ (46a)
$A^{(2)}_{\rm spin}({\bf k},\omega)$ $=$ ${\bar{B}_{2{\bf k}}\over\pi\bar{\omega}_{2{\bf k}}}[\delta(\omega-\bar{\omega}_{2{\bf k}})-\delta(\omega+\bar{\omega}_{2{\bf k}})],$ (46b)
respectively. However, during the calculation, we have found that the spectral weight in the component $A^{(1)}_{\rm spin}({\bf k},\omega)$ is several orders of magnitude greater than the corresponding spectral weight in the component $A^{(2)}_{\rm spin}({\bf k},\omega)$, leading to that the electron umklapp scattering is mainly mediated by the spin excitations from the component of $A^{(1)}_{\rm spin}({\bf k},\omega)$. On the other hand, the spin excitations with the full spin excitation dispersion $\bar{\omega}_{1{\bf k}}$ around the AF wave vector ${\bf k}_{\rm A}=[\pm\pi,\pm\pi]$ have the largest density of states. With these special properties of the spin excitations in the underdoped pseudogap phase, the full effective spin propagator $P({\bf k},{\bf p},{\bf k}^{\prime},\omega)$ in Eq. (37) can be reduced approximately as,
$P({\bf k},{\bf p},{\bf k}^{\prime},\omega)$$\approx$ $-{1\over N}\sum\limits_{\bf q}\left[{\varpi^{(1)}_{11}({\bf k},{\bf p},{\bf k}^{\prime},{\bf q})\over\omega^{2}-[\omega^{(1)}_{11{\bf p}{\bf q}}]^{2}}\right.$ (47)
-- $\left.{\varpi^{(2)}_{11}({\bf k},{\bf p},{\bf k}^{\prime},{\bf q})\over\omega^{2}-[\omega^{(2)}_{11{\bf p}{\bf q}}]^{2}}\right].~{}~{}~{}~{}~{}$
Now we follow the discussions of the nature of the kernel function in Ref. Ma23, in the overdoped strange-metal phase to make a taylor expansion for the spin excitation energy dispersions $\bar{\omega}^{(1)}_{11{\bf p}{\bf q}}=\bar{\omega}_{1{\bf q}+{\bf p}}+\bar{\omega}_{1{\bf q}}$ and $\bar{\omega}^{(2)}_{11{\bf p}{\bf q}}=\bar{\omega}_{1{\bf q}+{\bf p}}-\bar{\omega}_{1{\bf q}}$ in Eq. (34), and then the spin excitation energy dispersions $\bar{\omega}^{(1)}_{11{\bf p}{\bf q}}$ and $\bar{\omega}^{(2)}_{11{\bf p}{\bf q}}$ can be expressed approximately as,
$\bar{\omega}^{(1)}_{11{\bf p}{\bf q}}$ $=$ $\bar{\omega}_{1{\bf q}+{\bf p}}+\bar{\omega}_{1{\bf q}}\approx b_{\bf q}p^{2}+2\bar{\omega}_{1{\bf q}},$ (48a)
$\bar{\omega}^{(2)}_{11{\bf p}{\bf q}}$ $=$ $\bar{\omega}_{1{\bf q}+{\bf p}}-\bar{\omega}_{1{\bf q}}\approx b_{\bf q}p^{2},$ (48b)
with $b_{\bf q}=d^{2}\bar{\omega}_{1{\bf q}}/(d^{2}q)$. The above results in Eq. (48) show that as in the case of the the overdoped strange-metal phaseMa23 , the full effective spin propagator $P({\bf k},{\bf p},{\bf k}^{\prime},\omega)$ in Eq. (47) scales with $p^{2}$, and then when the electron umklapp scattering kicks in, the energy scale is proportional to $\Delta^{2}_{p}$ due to the presence of this $p^{2}$ scaling in the full effective spin propagator (47). In this case, $T_{\rm scale}=\bar{b}\Delta^{2}_{p}$ can be identified as the low-temperature scale,
Figure 9: The low-temperature scale $T_{\rm scale}$ as a function of doping.
with the average value $\bar{b}=(1/N)\sum\limits_{{\bf q}\in\{{\bf k}_{\rm A}\}}b({\bf q})$ that is a constant at a given doping, where the summation ${\bf q}\in\{{\bf k}_{\rm A}\}$ is restricted to the extremely small area $\{{\bf k}_{\rm A}\}$ around the ${\bf k}_{\rm A}$ point of BZ. However, this low-temperature scale $T_{\rm scale}$ is strongly doping dependent. To see this doping dependence of $T_{\rm scale}$ more clearly, we plot $T_{\rm scale}$ as a function of doping in Fig. 9, where $T_{\rm scale}$ is relatively high at around the slightly underdoped region, then it decreases when the doping is increased in the heavily underdoped region, and is reduced to a very low temperature in the overdoped regime. Moreover, $T_{\rm scale}$ as a function of doping in Fig. 9 in the underdoped regime presents a similar behavior of the antinodal spin pseudogap crossover temperature $T^{*}_{\rm s}$ shown in Fig. 2b, suggesting a possible correlation between $T_{\rm scale}$ and the opening of the antinodal spin pseudogap below $T^{*}_{\rm s}$ in the underdoped pseudogap phase.
With the help of the above full effective spin propagator $P({\bf k},{\bf p},{\bf k}^{\prime},\omega)$ in Eq. (47) and the doping dependence of $T_{\rm scale}$ in Fig. 9, we now follow the similar analysis carried out in the overdoped strange-metal phaseMa23 to show that two primary regions of the low-temperature resistivity need to be distinguished:
(i) in the low-temperature T-quadratic region ($T<T_{\rm scale}$) in the underdoped regime, the kernel function $F(\theta,\theta^{\prime})$ is reduced as $F(\theta,\theta^{\prime})\propto T^{2}$, which naturally produces a low-temperature T-quadratic resistivity $\rho(T)\propto T^{2}$ as shown in Fig. 5;
(ii) in the low-temperature T-linear region ($T>T_{\rm scale}$) in the overdoped regime, the kernel function $F(\theta,\theta^{\prime})$ is reduced as $F(\theta,\theta^{\prime})\propto T$, which naturally induces a T-linear resistivity $\rho(T)\propto T$ as we have shown in Ref. Ma23, . In particular, this low-temperature scale $T_{\rm scale}$ in the overdoped regime is very low Ma23 due to the absence of the antinodal spin pseudogap as shown in Fig. 2.
IV Summary and discussion
Within the framework of the kinetic-energy-driven superconductivity, we have rederived the full spin propagator, the full charge-carrier propagator, and the full electron propagator in the normal-state of cuprate superconductors, where the spin self-energy (then the spin pseudogap) in the normal-state is obtained explicitly in terms of the collective charge-carrier mode in the particle-hole channel, and the charge-carrier self-energy (then the charge-carrier pseudogap) in the normal-state is obtained explicitly in terms of the spin excitation mode, while the electron self-energy (then the normal-state pseudogap) originated from the charge-carrier self-energy (then the charge-carrier pseudogap) is due to the charge-spin recombination. Moreover, we have also shown that (i) the spin excitation energy dispersion is anisotropically renormalized due to the momentum dependence of the spin pseudogap. In particular, the antinodal spin pseudogap effect is particularly notable in the slightly underdoped region, and then this antinodal spin pseudogap rapidly decreases with the increase of doping in the heavily underdoped region, eventually abrupt disappearing at around the optimal doping; (ii) the electronic density of state at around the antinodal region is gapped out by the normal-state pseudogap, and then the closed EFS contour is truncated to a set of four disconnected Fermi arcs centered at around the nodal region. By virtue of this full spin propagator and the reconstructed EFS, we have investigated the low-temperature electrical transport in the underdoped pseudogap phase of cuprate superconductors, where the scattering rate originated from the umklapp scattering between electrons by the exchange of the full effective spin propagator is derived within the framework of the Boltzmann transport theory. Our results show that the dominant contribution to the low-temperature electrical transport mainly comes from the antinodal umklapp scattering. In particular, a low temperature $T_{\rm scale}$ scales with $\Delta^{2}_{p}$ in the underdoped regime due to the opening of the momentum dependence of the spin pseudogap, where $\Delta_{p}$ is the minimal umklapp vector at the antinode. Moreover, this $T_{\rm scale}$ as a function of doping presents a similar behavior of the antinodal spin pseudogap crossover temperature $T^{*}_{\rm s}$, i.e., $T_{\rm scale}$ decreases with the increase of doping in the underdoped regime, and then is reduced to a very low temperature in the overdoped regime, which suggests a possible correlation between $T_{\rm scale}$ and the opening of the antinodal spin pseudogap below $T^{*}_{\rm s}$ in the underdoped pseudogap phase. In the underdoped regime, the resistivity exhibts a T-quadratic behaviour in the low-temperature region below $T_{\rm scale}$, where the strength of the T-quadratic resistivity decreases with the increase of doping. However, in the overdoped regime, the resistivity is T-linear in the low-temperature region above $T_{\rm scale}$. The current findings together with the recent resultsMa23 of the low-temperature T-linear resistivity in the overdoped strange-metal phase therefore show that (i) the electron umklapp scattering from a spin excitation associated with the antinodes leads to the T-linear resistivity in the weak coupling overdoped strange-metal phase; (ii) as this electron umklapp scattering flows to the strong coupling underdoped regime, the opening of the momentum dependence of the spin pseudogap lowers the density of stats of the spin excitations at around the antinodal region in the response to the intense umklapp scattering, which reduces the strength of the umklapp scattering from electronic states into the antinodal region, and therefore leads to the low-temperature T-quadratic form of the umklapp scattering rate. As a natural consequence, the low-temperature T-linear resistivity in the overdoped strange-metal phase is dramatically switched to the low-temperature T-quadratic resistivity in the underdoped pseudogap phase.
Acknowledgements
This work is supported by the National Key Research and Development Program of China under Grant Nos. 2023YFA1406500 and 2021YFA1401803, and the National Natural Science Foundation of China (NSFC) under Grant Nos. 12274036 and 12247116. H.G. acknowledge support from NSFC grant Nos. 11774019 and 12074022.