### v2

# Low-temperature T2 resistivity in the underdoped pseudogap phase versus T-linear resistivity in the overdoped strange-metal phase of cuprate superconductors

Abstract
The transport experiments demonstrate a dramatic switch from the low-temperature linear in temperature (T-linear) resistivity in the overdoped strange-metal phase of cuprate superconductors to the low-temperature quadratic in temperature (T-quadratic) resistivity in the underdoped pseudogap phase, however, a consensus on the origin of this unusual switch is still lacking. Here the resistivity in the underdoped pseudogap phase of cuprate superconductors is investigated using the Boltzmann transport equation. The resistivity originates from the electron umklapp scattering mediated by the spin excitation, however, the dominant contribution mainly comes from the antinodal umklapp scattering. In particular, a low temperature $T_{\rm scale}$ scales with $\Delta^{2}_{p}$ in the underdoped regime due to the opening of a momentum dependent spin pseudogap, where $\Delta_{p}$ is the minimal umklapp vector at the antinode. Moreover, this $T_{\rm scale}$ decreases with the increase of doping in the underdoped regime, and then is reduced to a very low temperature in the overdoped regime. In the underdoped regime, the resistivity is T-quadratic at the low temperatures below $T_{\rm scale}$, where the strength of the T-quadratic resistivity weakens as the doping is raised. However, in the overdoped regime, the resistivity is T-linear at the low temperatures above $T_{\rm scale}$. The results in this paper together with the recent study on the resistivity in the overdoped regime therefore show that the electron umklapp scattering from a spin excitation responsible for the low-temperature T-linear resistivity in the overdoped regime naturally produces the low-temperature T-quadratic resistivity in the underdoped regime resulting from the opening of a momentum dependent spin pseudogap.
pacs:
74.25.Fy, 74.25.Nf, 74.20.Mn, 74.72.-h
I Introduction
It has become very clear that cuprate superconductors are among the most complicated systems studied in condensed matter physics Bednorz86 ; Damascelli03 ; Campuzano04 ; Fink07 ; Deutscher05 ; Devereaux07 ; Fischer07 . The complications stem mainly from a fact that the parent compound of cuprate superconductors is an antiferromagnetic (AF) insulator Bednorz86 ; Damascelli03 ; Campuzano04 ; Fink07 ; Deutscher05 ; Devereaux07 ; Fischer07 . Superconductivity then is achieved by chemically introducing charge carriers to this AF insulator, which leads to that the physical properties mainly depend on the extent of doping, and the regimes have been classified into the underdoped, optimally doped, and overdoped, respectively. After intensive studies about four decades, a substantial amount of reliable and reproducible data has been accumulated by using many probes Damascelli03 ; Campuzano04 ; Fink07 ; Deutscher05 ; Devereaux07 ; Fischer07 ; Hussey08 ; Kastner98 ; Timusk99 ; Hufner08 ; Vishik18 ; Fujita12 , which reveals that the most of the unusual features of cuprate superconductors are observed in the normal-state. In particular, below a pseudogap crossover temperature $T^{*}$, which can be well above the superconducting (SC) transition temperature $T_{\rm c}$ in the underdoped regime, the physical response can be well interpreted in terms of the formation of a pseudogap by which it means a large suppression of the electronic density of states on the electron Fermi surface (EFS) Damascelli03 ; Campuzano04 ; Fink07 ; Deutscher05 ; Devereaux07 ; Fischer07 ; Hussey08 ; Kastner98 ; Timusk99 ; Hufner08 ; Vishik18 ; Fujita12 . This is why in the underdoped regime, the phase above $T_{\rm c}$ but below $T^{*}$ is so-called as the pseudogap phase. The pseudogap in the underdoped regime was first discovered through Knight-shift of the magnetic susceptibility experiments Warren89 ; Alloul89 , where the pseudogap is some kind of the spin pseudogap, meaning a strong reduction in the magnetic susceptibility through the particle-hole correlation. Subsequently, the presence of the pseudogap was confirmed by a series of experimental measurements taken with a wide variety of techniques Walstedt90 ; Takigawa91 ; Loeser96 ; Norman98 ; Renner98 ; Kohsaka08 ; Puchkov96 . Moreover, these experiments also show clearly that the pseudogap exists in both the spin and charge channels Timusk99 ; Hufner08 ; Vishik18 ; Fujita12 ; Warren89 ; Alloul89 ; Walstedt90 ; Takigawa91 ; Loeser96 ; Norman98 ; Renner98 ; Kohsaka08 ; Puchkov96 . For example, angle-resolved photoemission spectroscopy (ARPES) and scanning tunneling spectroscopy measure the charge channel Loeser96 ; Norman98 , while nuclear magnetic resonance (NMR) and nuclear quadrupole resonance (NQR) detect the spin channelFujita12 ; Warren89 ; Alloul89 ; Walstedt90 ; Takigawa91 . On the other hand, in the overdoped regime, the normal-state is characterized by a number of the anomalous low-temperature properties Damascelli03 ; Campuzano04 ; Fink07 ; Deutscher05 ; Devereaux07 ; Fischer07 ; Hussey08 ; Kastner98 ; Timusk99 ; Hufner08 ; Vishik18 ; Fujita12 in the sense that they do not fit in with the conventional Fermi-liquid theory (FL) Schrieffer64 ; Abrikosov88 ; Mahan81 . This has led to the normal-state in the overdoped regime being refereed to as the strange-metal phaseKeimer15 .
Among the notable characteristics in the normal-state, the most iconic feature has to be the electrical transportHussey08 ; Kastner98 ; Timusk99 ; Hufner08 . In the overdoped regime, a series of transport measurements revealed a T-linear resistivityKeimer15 ; Hussey23 ; Legros19 ; Ayres21 ; Grisso21 , where (i) the linear temperature term often extends to low temperatures of a few kelvin; and (ii) the linear temperature term persists to high temperatures with the same slope all the way to the lowest temperature. In this paper, we shall not address the case at high temperature, but instead focus on the low-temperature T-linear resistivity, which is in a striking contrast to the conventional FL behaviour in the conventional superconductorsSchrieffer64 ; Abrikosov88 ; Mahan81 . However, in the underdoped regime, the early transport measurements indicated that the opening of the pseudogap in the spin excitation spectrum leads to an obvious deviation from the low-temperature T-linear behaviour of the resistivityGurvitch87 . In particular, the subsequent transport experiments observed a reduced resistivity below $T^{*}$, near the temperature where the spin pseudogap, as seen by NMR Knight-shift and spin-relaxation rate, openedBucher93 ; Ito93 ; Nakano94 ; Ando01 . Thus the low-temperature resistivity anomaly coincides roughly with $T^{*}$ and these experimental results suggest that the decrease in the low-temperature resistivity below $T^{*}$ is caused by the reduced electron scattering by the spin excitations resulting from the opening the spin pseudogapBucher93 ; Ito93 ; Nakano94 ; Ando01 . Lately, the low-temperature resistivity was confirmed experimentally in a quadratic temperature dependenceAndo04a ; Ando04 ; Lee05 ; Proust08 , as would be expected from the conventional FL theorySchrieffer64 ; Abrikosov88 ; Mahan81 , for a wide doping range in the underdoped regime. More specifically, the recent transport experiments demonstrate a clear and dramatic switch from the low-temperature T-linear resistivity in the overdoped regime to the purely T-quadratic resistivity in the underdoped regimeCooper09 ; Hussey11 ; Mirzaei13 ; Barisic13 ; Pelc20 , where the strength of the T-quadratic resistivity decreases with the increase of doping. In spite of the considerable variation in crystal structures and impurity-scattering effects among different families of cuprate superconductors, these experimental results Gurvitch87 ; Bucher93 ; Ito93 ; Nakano94 ; Ando01 ; Ando04a ; Ando04 ; Lee05 ; Proust08 ; Cooper09 ; Hussey11 ; Mirzaei13 ; Barisic13 ; Pelc20 also show that as the low-temperature T-linear resistivity in the overdoped regime Legros19 ; Ayres21 ; Grisso21 , the T-quadratic resistivity also is a universal feature in the underdoped pseudogap phase. In this case, some crucial questions are raised: (i) why the low-temperature resistivity exhibits the T-quadratic behaviour in the underdoped regime, with an exotic crossover to the T-linear behaviour in the overdoped regime? (ii) whether the pseudogap is correlated with this exotic crossover or not? (iii) is there a common electron scattering mediated by the same bosonic mode that is responsible for both the low-temperature T-quadratic resistivity in the underdoped regime and low-temperature T-linear resistivity in the overdoped regime?
The theoretical explanation for this exotic crossover along with the underlying scattering mechanism is clearly a big challenge. For a understanding of the nature of the low-temperature T-linear resistivity in the overdoped regime, several scenarios have been proposed Varma89 ; Varma16 ; Varma20 ; Damle97 ; Sachdev11 ; Zaanen04 ; Luca07 ; Zaanen19 ; Hartnoll22 ; Lee21 . However, these scenarios are just as diverse as the mechanism of superconductivity. In the marginal FL phenomenologyVarma89 ; Varma16 ; Varma20 , a single T-linear scattering rate is posited to account for the T-linear resistivity. In particular, the T-linear resistivity was interpreted as a consequence of the scale invariant physics near to the quantum critical point (QCP) Varma20 ; Damle97 ; Sachdev11 . With the close relation to the physics of QCP, the T-linear resistivity has been interpreted in terms of the Planckian dissipationZaanen04 ; Luca07 ; Zaanen19 ; Hartnoll22 . Moreover, it has been shown recently that the low-temperature T-linear resistivity originates from the umklapp scattering between electrons by the exchange of a critical boson propagator Lee21 . These studiesLee21 together with many othersHonerkamp01 ; Hartnoll12 thus suggest that the electron umklapp scattering is the origin of the low-temperature T-linear resistivity. On the other hand, it has been argued that if the electron scattering responsible for the low-temperature T-linear resistivity in the overdoped regime involves the electron scattering on the spin excitations in the underdoped regime Puchkov96 , then the spin pseudogap seen in NMR below $T^{*}$ would naturally account for a deviation from the low-temperature T-linear behaviour of the resistivityBucher93 ; Ito93 ; Nakano94 ; Ando01 . In particular, the mechanism for the resistivity in the underdoped regime has been proposedHussey03 ; Rice17 , where the electron umklapp scattering above $T^{*}$ leads to the T-linear resistivityRice17 . As the temperatures fall below $T^{*}$, the pseudogap opens, and then restricts the available umklapp scattering channels, leading to a crossover between the T-linear resistivity at the temperature above $T^{*}$ and T-quadratic form of the scattering from the Fermi pockets at the temperature below $T^{*}$. Moreover, the low-temperature resistivity has been studied based on the two-particle self-consistent approachBergeron11 , where it has been shown that the resistivity is linear at low temperatures in the overdoped regime, however, in the underdoped regime, the resistivity may display a T-quadratic behaviour at temperatures below $T^{*}$. However, up to now, the T-quadratic resistivity in the underdoped pseudogap phase and its connection with the pseudogap is still the subject of much study, and the microscopic origin of the clear switch from the low-temperature T-linear resistivity in the overdoped regime to the low-temperature T-quadratic resistivity in the underdoped regime is hotly debated.
In the very recent studyMa23 , we have studied the nature of the low-temperature T-linear resistivity in the overdoped regime. In our scenario, the scattering rate arises from the umklapp scattering between electrons by the exchange of the effective spin propagator, where the dominant contribution mainly comes from the antinodal umklapp scattering. This umklapp scattering rate scales linearly with temperature in the low temperatures, which then naturally generates a low-temperature T-linear resistivity. In this paper, we study the low-temperature resistivity in the underdoped pseudogap phase along with this line, where in a striking difference to the case in the overdoped regime, the spin excitation energy dispersion is anisotropically renormalized due to the opening of a momentum dependent spin pseudogap. In particular, the density of the spin excitations at around the antinodal is heavily reduced by the antinodal spin pseudogap. In this case, a low temperature $T_{\rm scale}$ scales with $\Delta^{2}_{p}$, where $\Delta_{p}$ is the minimal umklapp vector at the antinode. Notably, this low-temperature scale $T_{\rm scale}$ as a function of doping presents a similar behavior of the antinodal spin pseudogap crossover temperature, i.e., $T_{\rm scale}$ is dropped down with the enhancement of doping in the underdoped regime, and then is reduced to a very low temperature in the overdoped regime. In the underdoped regime, the resistivity is T-quadratic at the low temperatures below $T_{\rm scale}$, where the T-quadratic resistivity strength weakens as the doping is raised. However, in the overdoped regime, the resistivity is instead T-linear in the low temperatures above $T_{\rm scale}$. The present results combined with the recent resultsMa23 on the low-temperature T-linear resistivity in the overdoped regime therefore reveal that (i) the electron umklapp scattering from a spin excitation associated with the antinodes leads to the T-linear resistivity in the weak coupling overdoped regime; (ii) as this electron umklapp scattering flows to the strong coupling underdoped regime, the opening of the momentum dependent spin pseudogap lowers the density of the spin excitations at around the antinodal region, which reduces the strength of the intense umklapp scattering from the electronic states into the antinodal region, and therefore leads to the low-temperature T-quadratic form of the umklapp scattering rate.
This paper is organized as follows. In the next section, we first show that (i) the antinodal spin pseudogap decreases with the increase of doping in the underdoped regime, and then it abruptly vanishes at around the optimal doping; (ii) the electronic density of state at around the antinodal region is gapped out by the normal-state pseudogap, and then the closed EFS contour is truncated to a set of four disconnected Fermi arcs centered at around the nodal region. Following this reconstructed EFS, the scattering rate originated from the umklapp scattering between electrons by the exchange of the full effective spin propagator is derived using the Boltzmann transport equation. The discussions of the quantitative characteristics of the low-temperature resistivity in the underdoped pseudogap phase are given in Section III, where we show that the scattering rate in the whole doping regime is predominantly governed by the antinodal umklapp scattering. In Section IV, we close the presentation with a summary and discussion of the main results. In Appendix A, we present the details for the derivation of the full spin and full electron propagators.
II Methodology
II.1 Low-energy effective $t$-$J$ model
The common element in the layered crystal structure of cuprate superconductors is the square-lattice copper-oxide layerBednorz86 , and then the unusual features mainly come from the strongly correlated motion of the electrons in the copper-oxide layerCooper94 ; Nakamura93 ; Hou94 ; Takenaka94 . Quickly after the discovery of superconductivity in cuprate superconductors, AndersonAnderson87 recognized that the essential physics of the doped copper-oxide layer can be modeled with the low-energy effective $t$-$J$ model,
$H$ $=$ $-t\sum_{\langle l\hat{\eta}\rangle\sigma}C^{\dagger}_{l\sigma}C_{l+\hat{\eta}\sigma}+t^{\prime}\sum_{\langle l\hat{\tau}\rangle\sigma}C^{\dagger}_{l\sigma}C_{l+\hat{\tau}\sigma}+\mu\sum_{l\sigma}C^{\dagger}_{l\sigma}C_{l\sigma}$ (1)
$+$ $J\sum_{\langle l\hat{\eta}\rangle}{\bf S}_{l}\cdot{\bf S}_{l+\hat{\eta}},~{}~{}~{}~{}$
where $C^{\dagger}_{l\sigma}$ ($C_{l\sigma}$) is electron operator that creates (annihilates) an electron on site $l$ with spin $\sigma$, ${\bf S}_{l}$ is the spin operator with its components $S_{l}^{x}$, $S_{l}^{y}$, and $S_{l}^{z}$, and $\mu$ is the chemical potential. This $t$-$J$ model (1) describes a competition between the kinetic energy and magnetic energy, and the kinetic energy includes the nearest-neighbor (NN) hopping with the hopping integral $t$ and next NN hopping with the hopping integral $t^{\prime}$, while the magnetic energy is described by an AF Heisenberg term with the NN coupling $J$. $\langle l\hat{\eta}\rangle$ ($\langle l\hat{\tau}\rangle$) indicates that $l$ runs over all sites, and for each $l$, over its NN sites $\hat{\eta}$ (next NN sites $\hat{\tau}$). The parameters in this paper are chosen as $t/J=2.5$ and $t^{\prime}/t=0.3$ as the recent discussions Ma23 . Moreover, we use a notation in which the magnetic coupling $J$ and the lattice constant of the square lattice are set to the energy and length units, respectively. However, when necessary to compare with the experimental data, we take $J=100$meV, which is the typical value of cuprate superconductors.
The $t$-$J$ model (1) is supplemented by a crucial on-site local constraint $\sum_{\sigma}C^{\dagger}_{l\sigma}C_{l\sigma}\leq 1$ that the double occupancy of a site by two electrons of opposite spins is not allowed Anderson87 , while the strong electron correlation in the system manifests itself by this no double electron occupancy local constraint Yu92 ; Lee06 ; Edegger07 ; Spalek22 ; Zhang93 . However, the no double electron occupancy also leads to the difficulty for the studying of the $t$-$J$ model (1). For a proper treatment of this no double electron occupancy local constraint, the fermion-spin transformationFeng0494 ; Feng15 has been established, where the physics of no double electron occupancy is taken into account by representing the constrained electron as,
$C_{l\uparrow}=h^{\dagger}_{l\uparrow}S^{-}_{l},~{}~{}~{}~{}C_{l\downarrow}=h^{\dagger}_{l\downarrow}S^{+}_{l},$ (2)
where the $U(1)$ gauge invariant spin-raising (spin-lowering) operator $S^{+}_{l}$ ($S^{-}_{l}$) carries spin index of the constrained electron, and therefore the collective mode from this spin degree of freedom of the constrained electron is interpreted as the spin excitation responsible for the spin dynamics of the system, and the $U(1)$ gauge invariant spinful fermion operator $h^{\dagger}_{l\sigma}=e^{i\Phi_{l\sigma}}h^{\dagger}_{l}$ ($h_{l\sigma}=e^{-i\Phi_{l\sigma}}h_{l}$) creates (annihilates) a charge carrier on site $l$, and therefore represents the charge degree of freedom of the constrained electron together with some effects of spin configuration rearrangements due to the presence of the doped charge carrier itself, while the charge carrier and localized spin recombine to form the physical electron responsible for the electronic properties.
II.2 Pseudogap in charge and spin channels
In the past three decades, a series of experiments from NMR, NQR, and the inelastic neutron scattering (INS) measurements Fujita12 ; Birgeneau89 ; Fong95 ; Yamada98 ; Arai99 ; Bourges00 ; He01 ; Tranquada04 ; Bourges05 has provided an intrinsic connection between the electron pairing mechanism and spin excitations in cuprate superconductors, where a key question is whether the spin excitation can mediate electron pairing in analogy to the phonon-mediate pairing mechanism in the conventional superconductorsSchrieffer64 . Starting from the $t$-$J$ model in the fermion-spin representation, the kinetic-energy-driven superconductivityFeng15 ; Feng0306 ; Feng12 ; Feng15a has been developed, where the d-wave charge-carrier pairing state is generated by the charge-carrier interaction directly from the kinetic energy of the $t$-$J$ model by the exchange of the spin excitation. However, the d-wave electron pairs originated from this d-wave charge-carrier pairing state are due to the charge-spin recombinationFeng15a , and these electron pairs condensation reveals the d-wave SC-state. This kinetic-energy-driven SC mechanism is purely electronic without phonon, since the glue to hold the constrained electron pairs together is the spin excitation, the collective mode from the spin degree of freedom of the constrained electron itself. Moreover, this electron paring state mediated by the spin excitation in a way is in turn strongly influenced by the electron coherence, which leads to that the maximal $T_{\rm c}$ occurs at the optimal doping, and then decreases in both the underdoped and overdoped regimes. Following these previous discussionsFeng15 ; Feng0306 ; Feng12 ; Feng15a , the full charge-carrier propagator in the normal-state can be derived as [see Appendix A],
$g({\bf k},\omega)={1\over\omega-\xi_{\bf k}-\Sigma^{({\rm h})}_{\rm ph}({\bf k},\omega)},$ (3)
where $\xi_{\bf k}=4t\chi_{1}\gamma_{{\bf k}}-4t^{\prime}\chi_{2}\gamma_{{\bf k}}^{\prime}-\mu_{\rm h}$ is the mean-filed (MF) charge-carrier energy dispersion, with $\gamma_{\bf k}=({\rm cos}k_{x}+{\rm cos}k_{y})/2$, $\gamma_{\bf k}^{\prime}={\rm cos}k_{x}{\rm cos}k_{y}$, and the spin correlation functions $\chi_{1}=\langle S^{+}_{l}S^{-}_{l+\hat{\eta}}\rangle$ and $\chi_{2}=\langle S_{l}^{+}S_{l+\hat{\tau}}^{-}\rangle$, while the charge-carrier normal self-energy $\Sigma^{({\rm h})}_{\rm ph}({\bf k},\omega)$ is obtained explicitly in Eq. (43) in Appendix A.
In the framework of the kinetic-energy-driven superconductivity Feng15 ; Feng0306 ; Feng12 ; Feng15a , the charge-carrier pseudogap forms due to the strong coupling between charge and spin degrees of freedom of the constrained electron. To explore the nature of this charge-carrier pseudogap more clearly, we rewrite the charge-carrier normal self-energy in Eq. (3) as,
$\Sigma^{({\rm h})}_{\rm ph}(\bf k,\omega)$$\approx$ ${[\bar{\Delta}^{({\rm h})}_{\rm pg}({\bf k})]^{2}\over\omega-\xi_{0{\bf k}}},$ (4)
where $\xi_{0{\bf k}}=L^{({\rm h})}_{2}({\bf k})/L^{({\rm h})}_{1}({\bf k})$ is the energy spectrum of $\Sigma^{({\rm h})}_{\rm ph}(\bf k,\omega)$, and $\bar{\Delta}^{({\rm h})}_{\rm pg}({\bf k})=L^{({\rm h})}_{2}({\bf k})/\sqrt{L^{({\rm h})}_{1}({\bf k})}$ is refereed to as the momentum dependence of the charge-carrier pseudogap, since it plays a role of the suppression of the low-energy spectral weight of the charge-carrier excitation spectrum, with the functions $L^{({\rm h})}_{1}({\bf k})=-\Sigma^{({\rm h})}_{\rm pho}({\bf k},\omega=0)$ and $L^{({\rm h})}_{2}({\bf k})=-\Sigma^{({\rm h})}_{\rm ph}({\bf k},\omega=0)$, while $\Sigma^{({\rm h})}_{\rm ph}({\bf k},\omega=0)$ and the antisymmetric part $\Sigma^{({\rm h})}_{\rm pho}({\bf k},\omega)$ of the charge-carrier normal self-energy are obtained directly from $\Sigma^{({\rm h})}_{\rm ph}({\bf k},\omega)$ in Eq. (43).
Figure 1: (a) The charge-carrier pseudogap at temperature $T=0.002J$ and (b) the charge-carrier pseudogap crossover temperature as a function of doping taken from Ref. Feng12, .
Moreover, the sharp peaks appearing at low-temperature in $\Sigma^{({\rm h})}_{\rm ph}(\bf k,\omega)$ and the related quantities are actually a $\delta$-function, which are broadened by a small damping employed in the numerical calculation for a finite latticeBrinckmann01 ; Restrepo23 . As the same approach of the numerical calculation carried out in Ref. Ma23, , the calculation in this paper for $\Sigma^{({\rm h})}_{\rm ph}(\bf k,\omega)$ and the related quantities is performed numerically on a $160\times 160$ lattice in momentum space, where the infinitesimal $i0_{+}\rightarrow i\Gamma$ is replaced by a small damping $\Gamma=0.05J$.
For a convenience in the following discussions, the charge-carrier pseudogap Feng15 ; Feng12 $\bar{\Delta}^{({\rm h})}_{\rm pg}$ at temperature $T=0.002J$ as a function of doping is replotted in Fig. 1a, where the relatively large $\bar{\Delta}^{({\rm h})}_{\rm pg}$ appears in the underdoped regime, and then it weakens as the optimal doping is approached. However, a quite weak $\bar{\Delta}^{({\rm h})}_{\rm pg}$ is still present at around the optimal doping, but it disappears at the heavily overdoped region. Moreover, at a given doping concentration, this $\bar{\Delta}^{({\rm h})}_{\rm pg}$ is identified with a crossover with a charge-carrier pseudogap crossover temperature $T^{*}_{\rm h}$ rather than a phase transition. To see the evolution of $T^{*}_{\rm h}$ with doping more clearly, $T^{*}_{\rm h}$ as a function of doping is also replotted in Fig. 1b, where in conformity with the doping dependence of $\bar{\Delta}^{({\rm h})}_{\rm pg}$ in Fig. 1a, $T^{*}_{\rm h}$ is relatively high at the slight underdoping, and then it smoothly decreases with the increase of doping in the underdoped regime, eventually terminating at the heavily overdoped regionFeng12 ; Feng15 . More importantly, as the electron pairing state originated from the charge-carrier pairing state are due to the charge-spin recombination, the normal-state pseudogap state originated from the charge-carrier pseudogap state is also due to the charge-spin recombinationFeng15a , and then the anomalous properties associated with the formation of the normal-state pseudogap are explained in a natural way Damascelli03 ; Campuzano04 ; Fink07 ; Deutscher05 ; Devereaux07 ; Fischer07 ; Hussey08 ; Kastner98 ; Timusk99 ; Hufner08 ; Vishik18 ; Fujita12 . We will return to the discussion of the normal-state pseudogap towards next subsection II.3.
On the other hand, in the fermion-spin theoryFeng0494 ; Feng15 , the scattering of spins due to the charge-carrier fluctuation dominates the spin dynamics. In this case, the dynamical spin response of cuprate superconductors has been discussed from the SC-state to the normal-stateKuang15 ; Yuan01 ; Feng98 , where the full spin propagator in the normal-state has been derived as [see Appendix A],
$D({\bf k},\omega)$ $=$ ${B_{\bf k}\over\omega^{2}-\omega^{2}_{\bf k}-B_{\bf k}\Sigma^{({\rm s})}_{\rm ph}({\bf k},\omega)},~{}~{}~{}~{}~{}$ (5)
with the MF spin excitation energy dispersion $\omega_{\bf k}$ and the corresponding weight function $B_{\bf k}$ of the MF spin excitation spectrum that have been obtained explicitly in the previous worksFeng15 ; Kuang15 , while the spin self-energy in the normal-state obtained in terms of the collective charge-carrier mode in the particle-hole channel that is given in Eq. (A.2) in Appendix A.
For a better understanding of the nonconventional features of the spin pseudogap, the above spin self-energy in Eq. (5) can be also rewritten as,
$\Sigma^{({\rm s})}_{\rm ph}({\bf k},\omega)\approx{B_{\bf k}[\bar{\Delta}^{({\rm s})}_{\rm pg}({\bf k})]^{2}\over\omega^{2}-\omega^{2}_{0{\bf k}}},$ (6)
where $\omega^{2}_{0{\bf k}}=L^{({\rm s})}_{1{\bf k}}/L^{({\rm s})}_{2{\bf k}}$ is the energy spectrum of $\Sigma^{({\rm s})}_{\rm ph}({\bf k},\omega)$, while $[\bar{\Delta}^{({\rm s})}_{\rm pg}({\bf k})]^{2}=[L^{({\rm s})}_{1{\bf k}}]^{2}/B_{\bf k}L^{({\rm s})}_{2{\bf k}}$ is identified as being a region of the spin self-energy in which $\bar{\Delta}^{({\rm s})}_{\rm pg}({\bf k})$ anisotropically reduces the density of the spin excitation, and in this sense, $\bar{\Delta}^{({\rm s})}_{\rm pg}({\bf k})$ is refereed to as the spin pseudogap. The functions $L^{({\rm s})}_{1{\bf k}}$ and $L^{({\rm s})}_{2{\bf k}}$ is derived directly from the spin self-energy $\Sigma^{({\rm s})}_{\rm ph}({\bf k},\omega)$ in Eq. (A.2) as,
$L^{({\rm s})}_{1{\bf k}}$ $=$ $-{B_{\bf k}\over N^{2}}\sum_{\bf pq}\Omega^{({\rm s})}_{{\bf k}{\bf p}{\bf q}}{F^{\rm(s)}({\bf k},{\bf p},{\bf q})\over[\omega_{{\bf q}+{\bf k}}-(\bar{\xi}_{{\bf p}+{\bf q}}-\bar{\xi}_{{\bf p}})]^{2}},~{}~{}~{}~{}~{}$ (7a)
$L^{({\rm s})}_{2{\bf k}}$ $=$ $-{B_{\bf k}\over N^{2}}\sum_{\bf pq}\Omega^{({\rm s})}_{{\bf k}{\bf p}{\bf q}}{F^{\rm(s)}({\bf k},{\bf p},{\bf q})\over[\omega_{{\bf q}+{\bf k}}-(\bar{\xi}_{{\bf p}+{\bf q}}-\bar{\xi}_{{\bf p}})]^{4}},~{}~{}~{}~{}~{}~{}~{}~{}$ (7b)
respectively, where the renormalized charge-carrier energy dispersion $\bar{\xi}_{\bf k}=Z^{({\rm h})}_{\rm F}\xi_{\bf k}$, with the charge-carrier coherent weight $Z^{({\rm h})-1}_{\rm F}=1-{\rm Re}\Sigma^{({\rm h})}_{\rm pho}({\bf k},0)\mid_{{\bf k}=[\pi,0]}$, while the function $F^{\rm(s)}({\bf k},{\bf p},{\bf q})$ and vertex function $\Omega^{({\rm s})}_{{\bf k}{\bf p}{\bf q}}$ are presented in Appendix A. Substituting the above spin self-energy (6) into Eq. (5), the full spin propagator in the normal-state can be derived as,
$D({\bf k},\omega)={\bar{B}_{1{\bf k}}\over\omega^{2}-\bar{\omega}^{2}_{1{\bf k}}}+{\bar{B}_{2{\bf k}}\over\omega^{2}-\bar{\omega}^{2}_{2{\bf k}}}=\sum_{\alpha=1,2}{\bar{B}_{\alpha{\bf k}}\over\omega^{2}-\bar{\omega}^{2}_{\alpha{\bf k}}},~{}~{}~{}~{}~{}~{}$ (8)
with the renormalized spin excitation energy dispersions,
$\bar{\omega}^{2}_{1{\bf k}}$ $=$ ${1\over 2}\left[\omega^{2}_{\bf k}+\omega^{2}_{0{\bf k}}+\sqrt{(\omega^{2}_{\bf k}-\omega^{2}_{0{\bf k}})^{2}+4B^{2}_{\bf k}[\bar{\Delta}^{({\rm s})}_{\rm pg}({\bf k})]^{2}}\right],$
$\bar{\omega}^{2}_{2{\bf k}}$ $=$ ${1\over 2}\left[\omega^{2}_{\bf k}+\omega^{2}_{0{\bf k}}-\sqrt{(\omega^{2}_{\bf k}-\omega^{2}_{0{\bf k}})^{2}+4B^{2}_{\bf k}[\bar{\Delta}^{({\rm s})}_{\rm pg}({\bf k})]^{2}}\right],$
and the corresponding weight functions of the full spin excitation spectrum,
Figure 2: (a) The spin pseudogap $\bar{\Delta}^{({\rm s})}_{\rm pg}({\bf k}_{\rm AN})$ at the antinode for temperature $T=0.002J$ and (b) the corresponding antinodal spin-pseudogap crossover temperature $T^{*}_{\rm s}$ as a function of doping, where ${\bf k}_{\rm AN}$ is the wave vector at the antinode.
$\bar{B}_{1{\bf k}}$ $=$ ${1\over 2}B_{\bf k}\left({\omega^{2}_{\bf k}-\omega^{2}_{0{\bf k}}\over\sqrt{(\omega^{2}_{\bf k}-\omega^{2}_{0{\bf k}})^{2}+4B^{2}_{\bf k}[\bar{\Delta}^{({\rm s})}_{\rm pg}({\bf k})}]^{4}}+1\right),~{}~{}~{}~{}$ (10a)
$\bar{B}_{2{\bf k}}$ $=$ $-{1\over 2}B_{\bf k}\left({\omega^{2}_{\bf k}-\omega^{2}_{0{\bf k}}\over\sqrt{(\omega^{2}_{\bf k}-\omega^{2}_{0{\bf k}})^{2}+4B^{2}_{\bf k}[\bar{\Delta}^{({\rm s})}_{\rm pg}({\bf k})}]^{4}}-1\right).~{}~{}~{}~{}~{}~{}~{}~{}$ (10b)
It should be emphasized that the equation (6) is an identity only in the case of $\omega=0$, however, as in the case of the charge-carrier pseudogap Feng15 ; Feng12 , it is a proper approximation for the low-energy case of $\omega\neq 0$.
The momentum dependence of the spin pseudogap $\bar{\Delta}^{({\rm s})}_{\rm pg}({\bf k})$ in Eq. (6) evolves strongly with doping and temperature, as that in the charge-carrier pseudogap Feng15 ; Feng12 . On the other hand, weMa23 have shown that the umklapp scattering between electrons by the exchange of the MF effective spin propagator can give a consistent description of the low-temperature T-linear resistivity in the overdoped strange-metal phase, where the dominant contribution to the scattering rate mainly comes from the antinodal umklapp scattering. In this case, we focus on the exotic features of the doping and temperature dependent spin pseudogap $\bar{\Delta}^{({\rm s})}_{\rm pg}({\bf k})$ at around the antinodal region for a convenience in the following discussions of the low-temperature resistivity in the underdoped pseudogap phase. In Fig. 2a, we plot $\bar{\Delta}^{({\rm s})}_{\rm pg}({\bf k}_{\rm AN})$ at the antinode as a function of doping with temperature $T=0.002J$, where ${\bf k}_{\rm AN}$ is the wave vector at the antinode of the Brillouin zone (BZ). Apparently, this antinodal spin pseudogap is insensitive to the doping in the slightly underdoped region, and then it falls off rapidly as the doping is grown in the heavily underdoped region. To our big surprise, this antinodal spin pseudogap abruptly disappears at around the optimal doping. This antinodal spin pseudogap formation in the underdoped regime lowers the density of the spin excitations in response to the intense electron umklapp scattering from the spin excitations associated with the antinodes. However, in the overdoped regime, the main properties of the antinodal spin excitations can be well described by the MF spin propagator. Furthermore, for a given doping, this antinodal spin pseudogap vanishes when temperature reaches the antinodal spin-pseudogap crossover temperature $T^{*}_{\rm s}$. To see this doping dependence of $T^{*}_{\rm s}$ more clearly, we plot $T^{*}_{\rm s}$ as a function of doping in Fig. 2b, where in corresponding to the result of $\bar{\Delta}^{({\rm s})}_{\rm pg}({\bf k}_{\rm AN})$ in Fig. 2a, $T^{*}_{\rm s}$ is relatively high in the slightly underdoped region, and then it decreases rapidly when the doping is increased in the heavily underdoped region, eventually disappearing at around the optimal doping. It should be noted that the disappearance of the antinodal spin pseudogap at around the optimal doping does not indicate the existence of QCP at around the optimal doping. This follows from a basic fact that the experimental observations in the optimally doped and overdoped regimes show that although the strength of the T-linear resistivity gradually diminishes as a function of doping, the low-temperature T-linear resistivity retains a finite value up to the edge of the SC dome Keimer15 ; Hussey23 ; Legros19 ; Ayres21 ; Grisso21 , which can be described by the momentum relaxation due to the umklapp scattering between electrons by the exchange of the MF effective spin propagatorMa23 , however, it may be a challenging issue for the usual picture of a single QCP doping associated with the strange-metal phaseVarma20 ; Damle97 ; Sachdev11 .
Based on the full spin propagator (8), the dynamical spin response of cuprate superconductors in the normal-state has been investigatedKuang15 ; Yuan01 ; Feng98 , where the obtained results show the existence of damped but well-defined dispersive spin excitations in the whole doping phase diagram. In particular, the spectral weight of the spin excitation spectrum at around the antinodal region is strongly suppressed by the antinodal spin pseudogapKuang15 . Moreover, the low-energy spin fluctuation is dominated by the process from the mobile charge carriers, while the high-energy spin excitation on the other hand retains roughly constant energy as a function of doping, with spectral weight and dispersion relation comparable to those in the corresponding SC-state. All these results are qualitatively consistent with the experimental observations Fujita12 ; Birgeneau89 ; Fong95 ; Yamada98 ; Arai99 ; Bourges00 ; He01 ; Tranquada04 ; Bourges05 .
II.3 Electronic structure in the underdoped pseudogap state
For the understanding of the nature of the electronic structure in cuprate superconductors, it is needed to derive the electron propagator, which is characterized by the full charge-spin recombinationFeng15a . Following the previous discussionsFeng15a , the full electron propagator in the normal-state can be derived as [see Appendix A],
$G({\bf k},\omega)={1\over\omega-\varepsilon_{\bf k}-\Sigma_{\rm ph}({\bf k},\omega)},$ (11)
where $\varepsilon_{\bf k}=-4t\gamma_{\bf k}+4t^{\prime}\gamma_{\bf k}^{\prime}+\mu$ is the electron energy dispersion in the tight-binding approximation, while the electron normal self-energy $\Sigma_{\rm ph}({\bf k},\omega)$ is given in Eq. (51) in Appendix A.
The electron normal self-energy in Eq. (11) originated from the charge-carrier normal self-energy in Eq. (3) is due to the charge-spin recombination Feng15a , indicating that the normal-state pseudogap state originated from the charge-carrier pseudogap state is due to the charge-spin recombination. To see the nature of the normal-state pseudogap more clearly, the above electron normal self-energy in Eq. (11) can be reexpressed as,
$\Sigma_{\rm ph}({\bf k},\omega)$$\approx$ ${[\bar{\Delta}_{\rm PG}({\bf k})]^{2}\over\omega-\varepsilon_{0{\bf k}}},$ (12)
where $\varepsilon_{0{\bf k}}=L_{2{\bf k}}/L_{1{\bf k}}$ is the energy spectrum of $\Sigma_{\rm ph}({\bf k},\omega)$, and $\bar{\Delta}^{2}_{\rm PG}({\bf k})=L^{2}_{2{\bf k}}/L_{1{\bf k}}$ is the normal-state pseudogap, with the functions $L_{1}({\bf k})=-\Sigma_{\rm pho}({\bf k},\omega=0)$ and $L_{2}({\bf k})=-\Sigma_{\rm ph}({\bf k},\omega=0)$, while the antisymmetric part $\Sigma_{\rm pho}({\bf k},\omega)$ of the electron normal self-energy can be obtained directly from the electron normal self-energy $\Sigma_{\rm ph}({\bf k},\omega)$ in Eq. (51). This normal-state pseudogap $\bar{\Delta}_{\rm PG}({\bf k})$ is also identified as being a region of the electron normal self-energy in which the normal-state pseudogap anisotropically suppresses the electronic density of states on EFS. Since the normal-state pseudogap state is induced by the charge-carrier pseudogap state, the normal-state pseudogap $\bar{\Delta}_{\rm PG}({\bf k})$ [the normal-state pseudogap crossover temperature $T^{*}$] as a function of doping presents a similar behavior of the charge-carrier pseudogap $\bar{\Delta}^{({\rm h})}_{\rm pg}({\bf k})$ [the charge-carrier pseudogap crossover temperature $T^{*}_{\rm h}$]Feng15 ; Feng12 . This normal-state pseudogap crossover temperature $T^{*}$ is actually a crossover line below which a novel electronic state emerges, as exemplified by the presence of the Fermi arcs, the competing electronic orders, etc., and then the unconventional features of this novel electronic state can be well interpreted in terms of the formation of the normal-state pseudogap Damascelli03 ; Campuzano04 ; Fink07 ; Deutscher05 ; Devereaux07 ; Fischer07 ; Hussey08 ; Kastner98 ; Timusk99 ; Hufner08 ; Vishik18 .
The measured energy and momentum distribution curves can be depicted by the electron excitation spectrum Dessau91 ; Randeria95 ; Fedorov99 ; Campuzano99 ; DMou17 ; Bogdanov00 ; Kaminski01 ; Johnson01 ; Iwasawa08 ; Plumb13 ,
$I({\bf k},\omega)\propto n_{\rm F}(\omega)A({\bf k},\omega),$ (13)
with the electron spectral function,
$A({\bf k},\omega)={1\over\pi}{{\rm Im}\Sigma_{\rm ph}({\bf k},\omega)\over[\omega-\varepsilon_{\bf k}-{\rm Re}\Sigma_{\rm ph}({\bf k},\omega)]^{2}+[{\rm Im}\Sigma_{\rm ph}({\bf k},\omega)]^{2}},~{}~{}~{}$ (14)
where ${\rm Re}\Sigma_{\rm ph}({\bf k},\omega)$ and ${\rm Im}\Sigma_{\rm ph}({\bf k},\omega)$ are the real and imaginary parts of $\Sigma_{\rm ph}({\bf k},\omega)$, respectively. The electrons are renormalzed due to the electron scattering mediated by the spin excitation and then they acquire a finite lifetime.
The EFS topology is known to be crucial in the understanding of the unconventional superconductivity in cuprate superconductors Damascelli03 ; Campuzano04 ; Fink07 ; Deutscher05 ; Devereaux07 ; Fischer07 and the related anomalous normal-state properties Hussey08 ; Kastner98 ; Timusk99 ; Hufner08 ; Vishik18 . The ARPES measured EFS can be described theoretically by the intensity map of the electron excitation spectrum (13) at zero energy $\omega=0$, where the locations of the EFS continuous contour in momentum space is determined directly by the poles of the electron spectral function (14), $\varepsilon_{\bf k}+{\rm Re}\Sigma_{\rm ph}({\bf k},0)=\bar{\varepsilon}_{\bf k}=0$, with the renormalized electron energy dispersion $\bar{\varepsilon}_{\bf k}=Z_{\rm F}\varepsilon_{\bf k}$ and the single-particle coherent weight $Z^{-1}_{\rm F}=1-{\rm Re}\Sigma_{\rm pho}({\bf k},0)\mid_{{\bf k}=[\pi,0]}$. However, the strong redistribution of the spectral weight on EFS is mainly governed by the momentum dependence of the normal-state pseudogap $\bar{\Delta}_{\rm PG}({\bf k})$ [then the single-electron scattering rate $\Gamma_{\bf k}(\omega)={\rm Im}\Sigma_{\rm ph}({\bf k},\omega)=\pi[\bar{\Delta}_{\rm PG}({\bf k})]^{2}\delta(\omega+\varepsilon_{0{\bf k}})]$.
Figure 3: (Color online) Electron Fermi surface at $\delta=0.09$ with $T=0.002J$, where the Brillouin zone center has been shifted by [$\pi,\pi$], and AN, TFA, and ND denote the antinode, tip of the Fermi arc, and node, respectively.
For a convenience in the following discussions of the resistivity in the underdoped pseudogap phase, the EFS mapFeng16 ; Liu21 ; Cao21 ; Zeng22 at doping $\delta=0.09$ with temperature $T=0.002J$ is replotted in Fig. 3, where the BZ center has been shifted by [$\pi,\pi$], and AN, TFA, and ND indicate the antinode, tip of the Fermi arc, and node, respectively. It thus shows clearly that the antinodal region of EFS becomes partially gapped, leading to that EFS consists, not of a closed contour, but only of four disconnected Fermi arcs centered around the nodes, in qualitative agreement with the experimental results Loeser96 ; Norman98 ; Chatterjee06 ; Shi08 ; Sassa11 ; Kaminski15 ; Loret17 ; Loret18 ; He14 ; Comin14 ; Comin16 . However, the most of the spectral weight on the Fermi arcs locates at around the tips of the Fermi arcs, indicating that the electrons at around the tips of the Fermi arcs have a largest density of states Chatterjee06 ; Shi08 ; Sassa11 ; Kaminski15 ; Loret17 ; Loret18 ; He14 ; Comin14 ; Comin16 , and then charge order is driven by this EFS instability, with a characteristic wave vector corresponding to the tips of the Fermi arcsFeng16 ; Comin14 ; Comin16 .
The elementary excitations are parameterized by the electron spectral function (14), which has given a consistent description of the renormalization of the electrons in the underdoped pseudogap phaseFeng16 ; Liu21 ; Cao21 ; Zeng22 , where the obtained results show that the single-electron scattering rate has a well-pronounced peak structure at around the antinodal and nodal regions, which leads to the remarkable peak-dip-hump structure in the energy distribution curve. Moreover, the dispersion kink is induced by the inflection point in the single-electron scattering rate, while the spectral weight at around the dispersion kink is reduced highly by the corresponding peak in the real part of the electron normal self-energy. All these results are in qualitative agreement with the corresponding experimental results Dessau91 ; Randeria95 ; Fedorov99 ; Campuzano99 ; DMou17 ; Bogdanov00 ; Kaminski01 ; Johnson01 ; Iwasawa08 ; Plumb13 . More importantly, the anisotropic suppression of the electronic density of states on EFS by the normal-state pseudogap can affect the electrical transport in two ways Timusk99 : (i) through the reduction of the number of electron current-carrying states as a normal-state pseudogap forms; and (ii) since the electron current carriers are scattered by the spin excitation, through the reduction in the density of spin excitations.
II.4 Electrical transport due to umklapp scattering from a spin excitation
We now turn to discuss the low-temperature electrical transport in the underdoped pseudogap phase. Theoretically, the Boltzmann transport equation is the cornerstone for the discussions of the electrical transportAbrikosov88 ; Mahan81 , since the Boltzmann transport equation is valid in the case of either the existence of the well-defined quasiparticles or the treatment of the electron interaction mediated by different bosonic modes within the Eliashberg approach. This builds on the following pioneering works: (i) in the early days of the electrical transport research in the conventional superconductorsPrange64 , Prange and Kadanoff demonstrated that in an electron-phonon system, a set of transport equations can be derived in the Migdal’s approximation, where the electron interaction mediated by the phonon leads to the electron self-energy and vertex correction. In particular, this coupled set of transport equations for the electron and phonon distribution functions is correct even in the case of the absence of the well-defined quasiparticlesPrange64 . Nevertheless, one of the forms of the electrical transport equation,
$e{\bf E}\cdot\nabla_{\bf k}f({\bf k})=I_{\rm e-e},$ (15)
is identical to the electrical Boltzmann transport equation suggested by Landau for the case in which the quasiparticle is well-definedAbrikosov88 ; Mahan81 , with the charge of an electron $e$, the electron distribution function in a homogeneous system $f({\bf k},t)$, and the electron-electron collision term $I_{\rm e-e}$. For a convenience in the following discussions, the external magnetic field ${\bf H}$ has been dropped, and only an external electric field ${\bf E}$ is applied to the system; (ii) however, this Boltzmann transport equation (15) is not specific to the electron interaction mediated by the phonon in the conventional superconductors, and has been confirmed recently that it is also valid for the system with the electron interaction mediated by other bosonic excitations Lee21 ; Ma23 . With the advantage of their insightLee21 ; Ma23 ; Prange64 , the electrical transport due to the electron scattering mediated by various kinds of bosonic modes can be evaluated in a simple way even in the case of the break down of the quasiparticle picture.
Figure 4: (Color online) Schematic picture of the electron umklapp scattering process, where an electron on a electron Fermi surface (left) is scattered by its partner on the umklapp electron Fermi surface (right). The Fermi wave vector of the tips of the Fermi arcs ${\rm k}^{\rm TFA}_{\rm F}$ is the radius of the circular electron Fermi surface (red), and then an electron on this circular electron Fermi surface (left) parametrized by the Fermi angle $\theta$ is scattered to a point parametrized by the Fermi angle $\theta^{\prime}$ on the umklapp electron Fermi surface (right) by the spin excitation carrying momentum ${\rm p}(\theta,\theta^{\prime})$. $\Delta_{p}$ is the minimal umklapp vector at the antinode (the Fermi angle $\theta=0$).
In the following discussions, we study the the low-temperature resistivity in the underdoped pseudogap phase based on the Boltzmann transport equation (15). For the calculation of this Boltzmann transport equation (15), the linear perturbation from the equilibrium in terms of the distribution function has been introduced Lee21 ; Ma23 ; Prange64 ,
$f({\bf k})$ $=$ $n_{\rm F}({\bar{\varepsilon}_{\bf k}})-{dn_{\rm F}({\bar{\varepsilon}_{\bf k}})\over d{\bar{\varepsilon}_{\bf k}}}\tilde{\Phi}({\bf k}),$ (16)
with the fermion distribution functions $n_{\rm F}(\omega)$, and the local shift of the chemical potential at a given patch of EFS $\tilde{\Phi}({\bf k})$, which satisfies the antisymmetric relation $\tilde{\Phi}(-{\bf k})=-\tilde{\Phi}({\bf k})$. We substitute the above result in Eq. (16) into Eq. (15), and then linearize the Boltzmann equation (15) as,
$e{\bf v}_{\bf k}\cdot{\bf E}{dn_{\rm F}({\bar{\varepsilon}_{\bf k}})\over d{\bar{\varepsilon}_{\bf k}}}=I_{\rm e-e},$ (17)
where ${\bf v}_{\bf k}=\nabla_{\bf k}{\bar{\varepsilon}_{\bf k}}$ is the electron velocity.
The electron-electron collision term $I_{\rm e-e}$ in the Boltzmann equation is directly connected with the electrical scattering mechanism Abrikosov88 ; Mahan81 . Although there is no consensus on the origin of the electrical transport in the underdoped regime to date, it is widely believed that the opening of the spin pseudogap induces a deviation from the T-linear behaviour of the resistivityGurvitch87 ; Bucher93 ; Ito93 ; Nakano94 ; Ando01 . In particular, it has been shown clearly that the electron umklapp scattering is the origin of the low-temperature T-linear resistivity in the strange-metal phase Lee21 ; Honerkamp01 ; Hartnoll12 . Moreover, weMa23 have also shown very recently that the low-temperature T-linear resistivity in the overdoped strange-metal phase originates from the umklapp scattering between electrons by the exchange of the MF effective spin propagator. In the following discussions, we will show that when this electron umklapp scattering mediated by the spin excitation in the overdoped regimeMa23 flows to the underdoped regime, the opening of the momentum dependence of the spin pseudogap naturally leads to a low-temperature T-quadratic behaviour of the resistivity. To see the electron umklapp scattering process more clearly, we show a schematic picture of the electron umklapp scattering processLee21 ; Ma23 in Fig. 4, where an electron on a circular EFS (left) is scattered by its partner on the umklapp EFS (right). It should be noted that the intensity map of EFS in Fig. 4 is the same as the EFS map in Fig. 3, while the Fermi wave vector of the tips of the Fermi arcs ${\rm k}^{\rm TFA}_{\rm F}$ is the radius of the EFS circle (red). Moreover, this circle EFS (red) connects all tips of the Fermi arcs, and then the most of the electronic density of states is concentred on this circular EFS.
Following these recent discussionsLee21 ; Ma23 , we can derive the electron-electron collision $I_{\rm e-e}$ in Eq. (17) in the underdoped pseudogap phase as,
$I_{\rm e-e}$ $=$ ${1\over N^{2}}\sum_{{\bf k}^{\prime},{\bf p}}{2\over T}|P({\bf k},{\bf p},{\bf k}^{\prime},\bar{\varepsilon}_{\bf k}-\bar{\varepsilon}_{{\bf k}+{\bf p}+{\bf G}})|^{2}\{\tilde{\Phi}({\bf k})+\tilde{\Phi}({\bf k^{\prime}})-\tilde{\Phi}({\bf k}+{\bf p}+{\bf G})-\tilde{\Phi}({\bf k}^{\prime}-{\bf p})\}$ (18)
$\times$ $n_{\rm F}(\bar{\varepsilon}_{\bf k})n_{\rm F}(\bar{\varepsilon}_{{\bf k}^{\prime}})[1-n_{\rm F}(\bar{\varepsilon}_{{\bf k}+{\bf p}+{\bf G}})][1-n_{\rm F}(\bar{\varepsilon}_{{\bf k}^{\prime}-{\bf p}})]\delta(\bar{\varepsilon}_{\bf k}+\bar{\varepsilon}_{\bf k^{\prime}}-\bar{\varepsilon}_{{\bf k}+{\bf p}+{\bf G}}-\bar{\varepsilon}_{{\bf k}^{\prime}-{\bf p}}),$
where ${\bf G}$ represents a set of reciprocal lattice vectors. It should be emphasized that the above electron umklapp scattering (18) is described as a scattering between electrons by the exchange of the effective spin propagator,
$P({\bf k},{\bf p},{\bf k}^{\prime},\omega)$ $=$ ${1\over N}\sum_{\bf q}\Lambda_{{\bf p}+{\bf q}+{\bf k}}\Lambda_{{\bf q}+{\bf k}^{\prime}}\bar{\Pi}({\bf p},{\bf q},\omega),~{}~{}~{}$ (19)
rather than the scattering between electrons via the emission and absorption of the spin excitationLee21 ; Ma23 , where $\Lambda_{{\bf k}}=4t\gamma_{\bf k}-4t^{\prime}\gamma_{\bf k}^{\prime}$ is the bare vertex function.
In cuprate superconductors, a small density of charge carriers is sufficient to destroy the AF long-range order (AFLRO). However, the doped charge carriers and the coupled spins organize themselves in a cooperative way to enhance both the electrons mobility and the AF short-range order (AFSRO) correlation, and then the spin excitations in the spin liquid state with AFSRO appear to survive from the underdoped regime to the overdoped regimeFujita12 . Moreover, it has been shown that the coupling strength of the electrons with the spin excitations gradually weakens with the increase of doping from a strong-coupling case in the underdoped regime to a weak-coupling side in the overdoped regime Feng12 ; Johnson01 ; Kordyuk10 , reflecting a reduction of the strength of the magnetic fluctuation with the increase of doping. In other words, (i) the effect of the magnetic fluctuation in the overdoped regime is less dramatic than in the underdoped regimeKastner98 ; Timusk99 ; Hufner08 ; Vishik18 ; Fujita12 ; (ii) on the other hand, as shown in Fig. 2, the effect from antinodal spin pseudogap is absent in the overdoped regime, and then the main properties of the antinodal spin excitations can be well described by the MF spin propagator. These are reasons why the interpretation of the low-temperature T-linear resistivity in the overdoped strange-metal phase can be well made in terms of the umklapp scattering between electrons by the exchange of the MF effective spin propagatorMa23 .
However, in the strong-coupling side (then in the underdoped regime), (i) the effect of the magnetic fluctuation is much dramatic Kastner98 ; Timusk99 ; Hufner08 ; Vishik18 ; Fujita12 ; and (ii) as shown in Fig. 2, the effect from antinodal spin pseudogap is particularly notable, where the antinodal spin pseudogap lowers the density of the spin excitationsKuang15 in response to the intense electron umklapp scattering, which reduces the strength of the electron umklapp scattering from the spin excitation states into the antinodal region. In this case, the umklapp scattering between electrons should be mediated by the exchange of the full effective spin propagator for a proper description of the low-temperature resistivity in the underdoped pseudogap phase.
Now our aim is to obtain the full effective spin propagator. The full spin bubble $\bar{\Pi}({\bf p},{\bf q},\omega)$ in Eq. (19) is a convolution of two full spin propagators, and can be derived directly from the full spin propagator (8) as,
$\bar{\Pi}({\bf p},{\bf q},\omega)=-\sum_{\begin{subarray}{c}\alpha=1,2\\ \alpha^{\prime}=1,2\end{subarray}}{\bar{W}^{(1)}_{\alpha\alpha^{\prime}{\bf p}{\bf q}}\over\omega^{2}-[\bar{\omega}^{(1)}_{\alpha\alpha^{\prime}{\bf p}{\bf q}}]^{2}}+{\bar{W}^{(2)}_{\alpha\alpha^{\prime}{\bf p}{\bf q}}\over\omega^{2}-[\bar{\omega}^{(2)}_{\alpha\alpha^{\prime}{\bf p}{\bf q}}]^{2}},~{}~{}~{}~{}~{}~{}$ (20)
where the spin excitation energy dispersions $\bar{\omega}^{(1)}_{\alpha\alpha^{\prime}{\bf p}{\bf q}}$ and $\bar{\omega}^{(2)}_{\alpha\alpha^{\prime}{\bf p}{\bf q}}$ are given by,
$\bar{\omega}^{(1)}_{\alpha\alpha^{\prime}{\bf p}{\bf q}}$ $=$ $\bar{\omega}_{\alpha{\bf q}+{\bf p}}+\bar{\omega}_{\alpha^{\prime}{\bf q}},~{}~{}~{}~{}$ (21a)
$\bar{\omega}^{(2)}_{\alpha\alpha^{\prime}{\bf p}{\bf q}}$ $=$ $\bar{\omega}_{\alpha{\bf q}+{\bf p}}-\bar{\omega}_{\alpha^{\prime}{\bf q}},$ (21b)
respectively, and the corresponding functions,
$\bar{W}^{(1)}_{\alpha\alpha^{\prime}{\bf p}{\bf q}}$ $=$ ${\bar{B}_{\alpha^{\prime}{\bf q}}\bar{B}_{\alpha{\bf q}+{\bf p}}\over 2\bar{\omega}_{\alpha^{\prime}{\bf q}}\bar{\omega}_{\alpha{\bf q}+{\bf p}}}\bar{\omega}^{(1)}_{\alpha\alpha^{\prime}{\bf p}{\bf q}}$ (22)
$\times$ $[n_{\rm B}(\bar{\omega}_{\alpha{\bf q}+{\bf p}})+n_{\rm B}(\bar{\omega}_{\alpha^{\prime}{\bf q}})+1],~{}~{}~{}~{}~{}$
$\bar{W}^{(2)}_{\alpha\alpha^{\prime}{\bf p}{\bf q}}$ $=$ ${\bar{B}_{\alpha^{\prime}{\bf q}}\bar{B}_{\alpha{\bf q}+{\bf p}}\over 2\bar{\omega}_{\alpha^{\prime}{\bf q}}\bar{\omega}_{\alpha{\bf q}+{\bf p}}}\bar{\omega}^{(2)}_{\alpha\alpha^{\prime}{\bf p}{\bf q}}$ (23)
$\times$ $[n_{\rm B}(\bar{\omega}_{\alpha{\bf q}+{\bf p}})-n_{\rm B}(\bar{\omega}_{\alpha^{\prime}{\bf q}})],$
where $n_{\rm B}(\omega)$ is the boson distribution function. From the above spin bubble (20), the full effective spin propagator (19) now can be derived as,
$P({\bf k},{\bf p},{\bf k}^{\prime},\omega)$ $=$ $-{1\over N}\sum\limits_{\alpha\alpha^{\prime}{\bf q}}\left[{\varpi^{(1)}_{\alpha\alpha^{\prime}}({\bf k},{\bf p},{\bf k}^{\prime},{\bf q})\over\omega^{2}-[\bar{\omega}^{(1)}_{\alpha\alpha^{\prime}{\bf p}{\bf q}}]^{2}}\right.$ (24)
-- $\left.{\varpi^{(2)}_{\alpha\alpha^{\prime}}({\bf k},{\bf p},{\bf k}^{\prime},{\bf q})\over\omega^{2}-[\bar{\omega}^{(2)}_{\alpha\alpha^{\prime}{\bf p}{\bf q}}]^{2}}\right],~{}~{}~{}~{}~{}$
with the weight functions,
$\varpi^{(1)}_{\alpha\alpha^{\prime}}({\bf k},{\bf p},{\bf k}^{\prime},{\bf q})$ $=$ $\Lambda_{{\bf k}+{\bf p}+{\bf q}}\Lambda_{{\bf q}+{\bf k}^{\prime}}\bar{W}^{(1)}_{\alpha\alpha^{\prime}{\bf p}{\bf q}},$ (25a)
$\varpi^{(2)}_{\alpha\alpha^{\prime}}({\bf k},{\bf p},{\bf k}^{\prime},{\bf q})$ $=$ $\Lambda_{{\bf k}+{\bf p}+{\bf q}}\Lambda_{{\bf q}+{\bf k}^{\prime}}\bar{W}^{(2)}_{\alpha\alpha^{\prime}{\bf p}{\bf q}}.$ (25b)
The electron umklapp scattering in Eq. (18) shows that the electron-electron collision $I_{\rm e-e}$ is both functions of momentum and energy. Howerver, at low temperatures, everything happens near EFSAbrikosov88 ; Mahan81 . In this case, an any given patch on the circular EFS shown in Fig. 4 is represented via the Fermi angle $\theta$ with the Fermi angle range $\theta\in[0,2\pi]$, and then the momentum integration along the perpendicular momentum can be replaced by the integrationLee21 ; Ma23 ; Prange64 over $\bar{\varepsilon}_{\bf k}$. For the umklapp scattering between electrons by the exchange of the full effective spin propagator in Eq. (18), an electron on the circular EFS parametrized by the Fermi angle $\theta$ is scattered to a point parametrized by the Fermi angle $\theta^{\prime}$ on the umklapp EFS via the spin excitation carrying momentum ${\rm p}(\theta,\theta^{\prime})$ as shown in Fig. 4. Following the above treatment and the calculation process in the recent worksLee21 ; Ma23 for the low-temperature resistivity in the overdoped strange-metal phase, the electron-electron collision $I_{\rm e-e}$ in Eq. (18) in the underdoped pseudogap phase can be obtained straightforwardlyMa23 , and then the Boltzmann transport equation (17) can be derived as,
$e{\bf v}_{\rm F}(\theta)\cdot{\bf E}=-2\int{d\theta^{\prime}\over{2\pi}}\zeta(\theta^{\prime})F(\theta,\theta^{\prime})[\Phi(\theta)-\Phi(\theta^{\prime})],~{}~{}~{}~{}~{}$ (26)
with $\Phi(\theta)=\tilde{\Phi}[{\rm k}(\theta)]$, the Fermi velocity ${\bf v}_{\rm F}(\theta)$ at the Fermi angle $\theta$, the density of states factor $\zeta(\theta^{\prime})={\rm k}^{2}_{\rm F}/[4\pi^{2}{\rm v}^{3}_{\rm F}]$ at angle $\theta^{\prime}$, the Fermi wave vector ${\rm k}_{\rm F}$, and the Fermi velocity ${\rm v}_{\rm F}$. In particular, the antisymmetric relation $\tilde{\Phi}(-{\bf k})=-\tilde{\Phi}({\bf k})$ for $\tilde{\Phi}({\bf k})$ in Eq. (16) has been converted into $\Phi(\theta)=-\Phi(\theta+\pi)$ for $\Phi(\theta)$ in the above Eq. (26). Moreover, the coefficient of $\Phi(\theta)$ in the first term of the right-hand side of Eq. (26),
$\gamma(\theta)=2\int{d\theta^{\prime}\over{2\pi}}\zeta(\theta^{\prime})F(\theta,\theta^{\prime}),$ (27)
can be referred to as the angular (momentum) dependence of the umklapp scattering rateMa23 ; Lee21 , with the kernel function $F(\theta,\theta^{\prime})$ that connects the point $\theta$ on the circular EFS with the point $\theta^{\prime}$ on the umklapp EFS via the amplitude of the momentum transfer ${\rm p}(\theta,\theta^{\prime})$ as shown in Fig. 4, which can be derived as,
$F(\theta,\theta^{\prime})$ $=$ ${1\over T}\int{d\omega\over 2\pi}{\omega^{2}\over{\rm p}(\theta,\theta^{\prime})}{|\bar{P}[{\rm k}(\theta),{\rm p}(\theta,\theta^{\prime}),\omega]|}^{2}$ (28)
$\times$ $n_{\rm B}(\omega)[1+n_{\rm B}(\omega)],~{}~{}~{}~{}~{}~{}$
where the full effective spin propagator $\bar{P}[{\rm k}(\theta),{\rm p}(\theta,\theta^{\prime}),\omega]$ has been rewritten explicitly in terms of the Fermi angles $\theta$ and $\theta^{\prime}$ asMa23 ,
$\bar{P}[{\rm k}(\theta),{\rm p}(\theta,\theta^{\prime},\omega]$ $=$ $-{1\over N}\sum\limits_{\alpha\alpha^{\prime}{\bf q}}\left[{\varpi^{(1)}_{\alpha\alpha^{\prime}}(\theta,\theta^{\prime},{\bf q})\over\omega^{2}-[\bar{\omega}^{(1)}_{\alpha\alpha^{\prime}\theta,\theta^{\prime}}({\bf q})]^{2}}\right.$ (29)
-- $\left.{\varpi^{(2)}_{\alpha\alpha^{\prime}}(\theta,\theta^{\prime},{\bf q})\over\omega^{2}-[\bar{\omega}^{(2)}_{\alpha\alpha^{\prime}\theta,\theta^{\prime}}({\bf q})]^{2}}\right],~{}~{}~{}~{}~{}$
where $\varpi^{(1)}_{\alpha\alpha^{\prime}}(\theta,\theta^{\prime},{\bf q})=\varpi^{(1)}_{\alpha\alpha^{\prime}}[{\rm k}(\theta),{\rm p}(\theta,\theta^{\prime}),{\bf k}^{\prime}_{\rm F},{\bf q}]$, $\varpi^{(2)}_{\alpha\alpha^{\prime}}(\theta,\theta^{\prime},{\bf q})=\varpi^{(2)}_{\alpha\alpha^{\prime}}[{\rm k}(\theta),{\rm p}(\theta,\theta^{\prime}),{\bf k}^{\prime}_{\rm F},{\bf q}]$, $\bar{\omega}^{(1)}_{\alpha\alpha^{\prime}\theta,\theta^{\prime}}({\bf q})=\bar{\omega}^{(1)}_{\alpha\alpha^{\prime}{\rm p}(\theta,\theta^{\prime}){\bf q}}$, and $\bar{\omega}^{(2)}_{\alpha\alpha^{\prime}\theta,\theta^{\prime}}({\bf q})=\bar{\omega}^{(2)}_{\alpha\alpha^{\prime}{\rm p}(\theta,\theta^{\prime}){\bf q}}$.
At low temperatures, the electron elastic scattering occurs on EFS, while the electron inelastic scattering occurs near EFSKohsaka08 ; Abrikosov88 ; Mahan81 . In this case, the above energy integration in the kernel function (28) actually includes the inelastic scattering process of momentum and energy exchange between electronsLee21 ; Ma23 ; Prange64 . This follows from a fact that the umklapp scattering rate $\gamma(\theta)$ in Eq. (27) is directly associated with the kernel function $F(\theta,\theta^{\prime})$ in Eq. (28), while this kernel function is obtained by the integration of the spin excitation energy $\omega$, leading to the occurrence of the inelastic scattering process near EFS in the present umklapp scattering. In particular, in the case of the absence of the spin pseudogap, this electron umklapp scattering induces a low-temperature T-linear resistivity in the overdoped regimeMa23 , in agreement with the experimental observations Legros19 ; Ayres21 ; Grisso21 , where the experimental analyses seem to indicate that the T-linear resistivity is due to the inelastic scattering.
For the discussions of the low-temperature resistivity, we need to obtain the electron current density, which can be derived in terms of the local shift of the chemical potential $\Phi(\theta)$ asMa23 ,
${\bf J}$ $=$ $en_{0}{1\over N}\sum_{\bf k}{\bf v}_{\bf k}{dn_{\rm F}({\bar{\varepsilon}_{\bf k}})\over d\bar{\varepsilon}_{\bf k}}\tilde{\Phi}({\bf k})$ (30)
$=$ $-en_{0}{{\rm k}_{\rm F}\over{\rm v}_{\rm F}}\int{d\theta\over(2\pi)^{2}}{\bf v}_{\rm F}(\theta)\Phi(\theta),~{}~{}~{}~{}~{}$
with the momentum relaxation that is generated by the action of the electric field on the mobile electrons at EFS with the density $n_{0}$. However, this local shift of the chemical potential $\Phi(\theta)$ can be evaluated directly in the relaxation-time approximation asLee21 ; Ma23 , $\Phi(\theta)=-e{\rm v}_{\rm F}{\rm cos}(\theta)E_{\hat{x}}/[2\gamma(\theta)]$, where the electric field ${\bf E}$ has been selected along the $\hat{x}$-axis. With the help of the above treatment, the dc conductivity can be derived as Lee21 ; Ma23 ,
$\sigma_{\rm dc}(T)={1\over 2}e^{2}n_{0}{\rm k}_{\rm F}{\rm v}_{\rm F}\int{d\theta\over(2\pi)^{2}}{\rm cos}^{2}(\theta){1\over\gamma(\theta)},$ (31)
and then the resistivity can be expressed directly in terms of the above dc conductivity as,
$\rho(T)={1\over\sigma_{\rm dc}(T)}.$ (32)
It thus shows that the electrical resistance originates from the umklapp scattering between electrons by the exchange of the full effective spin propagator.
III Quantitative characteristics
Figure 5: (Color online) (a) Resistivity as a function of temperature at $\delta=0.06$ (black-line), $\delta=0.09$ (red-line), and $\delta=0.12$ (blue-line). Inset: the numerical fit of the resistivity (black-dots) at $\delta=0.09$ with the fit form $\rho(T)=A_{2}T^{2}$. (b) The corresponding experimental result of the low-temperature T-quadratic resistivity for the underdoped HgBa2CuO4+δ taken from Ref. Mirzaei13, .
It should be emphasized that the above resistivity in Eq. (32) is obtained in the pure two-dimensional $t$-$J$ model (1) on a square lattice. However, for a clear comparison with the corresponding experimental data Ando04a ; Ando04 ; Lee05 ; Proust08 ; Cooper09 ; Hussey11 ; Mirzaei13 ; Barisic13 ; Pelc20 , the resistivity obtained in the above equation (32) should be renormalized by the distance between the adjacent copper-oxide layers as it has been done in Ref. Bergeron11, , with the interlayer lattice constant Bergeron11 that can be chosen as $d=0.5$nm. Now we are ready to discuss the low-temperature resistivity in the underdoped pseudogap phase. In Fig. 5a, we plot the renormalized resistivity $\rho(T)$ as a function of temperature at doping $\delta=0.06$ (black-line), $\delta=0.09$ (red-line), and $\delta=0.12$ (blue-line), where in the low-temperature region, a power-law resistivity appears over a wide doping range in the underdoped regime. To explore this power-law behaviour of the low-temperature resistivity more clearly, the above results of $\rho(T)$ in Fig. 5a have been numerically fitted with the fit form $\rho(T)=A_{2}T^{2}$, and the fitted result of the low-temperature resistivity (black-dots) at $\delta=0.09$ is also plotted in Fig. 5a (inset), where the low-temperature resistivity is demonstrated clearly to be grown quadratically as the temperature is raised. For a better comparison, the corresponding experimental resultMirzaei13 of the low-temperature T-quadratic resistivity observed in the underdoped HgBa2CuO4+δ is also shown in Fig. 5b, where the most characteristic feature of the low-temperature resistivity in the underdoped regime is that it is perfectly quadratic down to the lowest achievable temperatures Ando04a ; Ando04 ; Lee05 ; Proust08 ; Cooper09 ; Hussey11 ; Mirzaei13 ; Barisic13 ; Pelc20 . Apparently, (i) this characteristic feature of the T-quadratic behaviour of the low-temperature resistivity is the same in the theory and experiments Ando04a ; Ando04 ; Lee05 ; Proust08 ; Cooper09 ; Hussey11 ; Mirzaei13 ; Barisic13 ; Pelc20 ; (ii) the magnitude of the low-temperature T-quadratic resistivity at a given doping and a given temperature is also qualitatively consistent with the corresponding experimental results in the underdoped regime Proust08 ; Cooper09 ; Hussey11 ; Mirzaei13 ; Barisic13 ; Pelc20 , where different magnitudes at a given doping and a given temperature have been observed for different families of cuprate superconductors.
Figure 6: (a) Strength of the T-quadratic resistivity as a function of doping. (b) The corresponding experimental result of cuprate superconductors taken from Ref. Barisic13, .
Moreover, the strength of the T-quadratic resistivity (then the T-quadratic resistivity coefficient) $A_{2}$ grows as the doping is reduced. To see this doping dependence of the T-quadratic resistivity strength more clearly, we plot $A_{2}$ as a function of doping $\delta$ in Fig. 6a in comparison with the corresponding experimental resultsBarisic13 of the underdoped cuprate superconductors in Fig. 6b. It thus shows clearly that $A_{2}$ drops gradually as doping is enhanced in the underdoped regime, indicating that $A_{2}$ is roughly proportional to the inverse of the doping concentration. This tendency of the doping dependent $A_{2}$ is also in qualitative agreement with the corresponding experimental results of cuprate superconductors in the underdoped regime Ando04a ; Ando04 ; Lee05 ; Proust08 ; Cooper09 ; Hussey11 ; Mirzaei13 ; Barisic13 ; Pelc20 . These results in Fig. 5 and Fig. 6 in the underdoped regime along with the recent resultsMa23 of the low-temperature T-linear resistivity in the overdoped regime therefore show that the electron umklapp scattering from a spin excitation responsible for the low-temperature T-linear resistivity in the overdoped regime naturally produces the low-temperature T-quadratic resistivity in the underdoped regime due to the opening of the momentum dependence of the spin pseudogap.
The low-temperature resistivity is mainly determined by the umklapp scattering rate $\gamma(\theta,T)$ in Eq. (27), however, this umklapp scattering rate $\gamma(\theta,T)$ is in turn directly connected with the kernel function $F(\theta,\theta^{\prime})$ (then the probability weight of the umklapp scattering) in Eq. (28) as we have mentioned above. For the further understanding of the nature of the umklapp scattering between electrons by the exchange of the full effective spin propagator, we first analyse the exotic feature of the kernel function $F(\theta,\theta^{\prime})$.
Figure 7: (Color online) Surface plot of the kernel function $F(\theta,\theta^{\prime})/F(\theta_{\rm AN},\theta^{\prime}_{\rm AN})$ at $\delta=0.09$ with $T=0.05J$, where AN, TFA, and ND denote the antinode, tip of the Fermi arc, and node, respectively, while $F(\theta_{\rm AN},\theta^{\prime}_{\rm AN})$ is the magnitude of $F(\theta,\theta^{\prime})$ at the antinode.
Figure 8: Umklapp scattering rate $\gamma(T)/\Delta\gamma$ at the antinode as a function of temperature for $\delta=0.09$, where $\Delta\gamma=\gamma(T_{0})-\gamma(0)$, with $T_{0}=0.13J$. The inset shows the detail of the temperature dependence of $\gamma(T)$ at the antinode in the far lower temperature region.
In Fig. 7, we plot $F(\theta,\theta^{\prime})$ at $\delta=0.09$ with $T=0.05J$, where the probability weight of the electron umklapp scattering is very strongly anisotropic in momentum space. In particular, as in the case of the overdoped strange-metal phaseMa23 , the almost all of the probability weight of the electron umklapp scattering in the underdoped pseudogap phase is also concentrated at around the antinodal region, leading to that the strongest umklapp scattering occurs at around this antinodal region. However, the weakest umklapp scattering appears at around the tips of the Fermi arcs, since the probability weight of the electron umklapp scattering there is almost none. Moreover, the strength of the electron umklapp scattering at around the nodal region is much weaker than that at around the antinodal region, since a very small amount of the probability weight of the electron umklapp scattering is concentrated at around the nodal region. This special momentum dependent distribution of the probability weight of the electron umklapp scattering in Fig. 7 together with the recent discussions of the momentum-dependent distribution of the probability weight of the electron umklapp scattering in the overdoped regimeMa23 therefore confirm that in the whole doping regime, the dominant contribution to the scattering rate mainly arises from the umklapp scattering between the relatively slow electrons at around the antinodal region.
We now turn to explore the unusual evolution of the umklapp scattering rate (27) with temperature in the underdoped pseudogap phase. We have made a series of calculations for the umklapp scattering rate $\gamma(T)$ at different Fermi angles, and the result of $\gamma(T)$ as a function of temperature for $\delta=0.09$ at the antinode is plotted in Fig. 8, where the inset shows the detail of the evolution of $\gamma(T)$ with temperature at the antinode in the far lower temperature region. Our numerical fit demonstrates that $\gamma(T)$ is purely T-quadratic in the low-temperature region, i.e., it increases quadratically with temperature as the temperature is increased. In particular, as the case of the resistivity shown in the inset of Fig. 5a, this $\gamma(T)$ approaches zero as the temperature approaches zero. Moreover, although the magnitude of $\gamma(\theta,T)$ is strongly anisotropic in momentum space, the low-temperature T-quadratic behaviour of $\gamma(\theta,T)$ appears at an any given Fermi angle $\theta$. Comparing this result in Fig. 8 with the corresponding results of the low-temperature T-quadratic resistivity in Fig. 5a, it thus shows that the low-temperature T-quadratic behaviour of $\gamma(T)$ together with the temperature region are the same as the corresponding behaviour and region in the resistivity $\rho(T)$, which therefore confirms that the low-temperature T-quadratic resistivity with the corresponding temperature region is mainly governed by the low-temperature T-quadratic umklapp scattering rate with the corresponding temperature region.
Now we give an explanation to show why the low-temperature resistivity has a T-quadratic behaviour in the underdoped pseudogap phase, with a dramatic switch to the T-linear behaviour in the overdoped strange-metal phase? The nature of the umklapp scattering rate in Eq. (27) is mainly determined by the nature of the kernel function $F(\theta,\theta^{\prime})$ in Eq. (28), however, this kernel function $F(\theta,\theta^{\prime})$ in the underdoped pseudogap phase is proportional to the full effective spin propagator $P({\bf k},{\bf p},{\bf k}^{\prime},\omega)$ in Eq. (19). From the full spin propagator in Eq. (8), the full spin spectral function is derived straightforwardly as $A_{\rm spin}({\bf k},\omega)=A^{(1)}_{\rm spin}({\bf k},\omega)+A^{(2)}_{\rm spin}({\bf k},\omega)$, where the components $A^{(1)}_{\rm spin}({\bf k},\omega)$ and $A^{(2)}_{\rm spin}({\bf k},\omega)$ are given by,
$A^{(1)}_{\rm spin}({\bf k},\omega)$ $=$ ${\bar{B}_{1{\bf k}}\over\pi\bar{\omega}_{1{\bf k}}}[\delta(\omega-\bar{\omega}_{1{\bf k}})-\delta(\omega+\bar{\omega}_{1{\bf k}})],~{}~{}~{}~{}$ (33a)
$A^{(2)}_{\rm spin}({\bf k},\omega)$ $=$ ${\bar{B}_{2{\bf k}}\over\pi\bar{\omega}_{2{\bf k}}}[\delta(\omega-\bar{\omega}_{2{\bf k}})-\delta(\omega+\bar{\omega}_{2{\bf k}})],$ (33b)
respectively. However, during the calculation, we have found that the spectral weight in the component $A^{(1)}_{\rm spin}({\bf k},\omega)$ is several orders of magnitude greater than the corresponding spectral weight in the component $A^{(2)}_{\rm spin}({\bf k},\omega)$, leading to that the electron umklapp scattering is mainly mediated by the spin excitations from the component of $A^{(1)}_{\rm spin}({\bf k},\omega)$.
Figure 9: The low-temperature scale $T_{\rm scale}$ as a function of doping.
On the other hand, the spin excitations with the full spin excitation dispersion $\bar{\omega}_{1{\bf k}}$ around the AF wave vector ${\bf k}_{\rm A}=[\pm\pi,\pm\pi]$ have the largest density of states. With these special properties of the spin excitations in the underdoped pseudogap phase, the full effective spin propagator $P({\bf k},{\bf p},{\bf k}^{\prime},\omega)$ in Eq. (24) can be reduced approximately as,
$P({\bf k},{\bf p},{\bf k}^{\prime},\omega)$$\approx$ $-{1\over N}\sum\limits_{\bf q}\left[{\varpi^{(1)}_{11}({\bf k},{\bf p},{\bf k}^{\prime},{\bf q})\over\omega^{2}-[\omega^{(1)}_{11{\bf p}{\bf q}}]^{2}}\right.$ (34)
-- $\left.{\varpi^{(2)}_{11}({\bf k},{\bf p},{\bf k}^{\prime},{\bf q})\over\omega^{2}-[\omega^{(2)}_{11{\bf p}{\bf q}}]^{2}}\right].~{}~{}~{}~{}~{}$
Now we follow the discussions of the nature of the kernel function in Ref. Ma23, in the overdoped strange-metal phase to make a taylor expansion for the spin excitation energy dispersions $\bar{\omega}^{(1)}_{11{\bf p}{\bf q}}=\bar{\omega}_{1{\bf q}+{\bf p}}+\bar{\omega}_{1{\bf q}}$ and $\bar{\omega}^{(2)}_{11{\bf p}{\bf q}}=\bar{\omega}_{1{\bf q}+{\bf p}}-\bar{\omega}_{1{\bf q}}$ in Eq. (21), and then the spin excitation energy dispersions $\bar{\omega}^{(1)}_{11{\bf p}{\bf q}}$ and $\bar{\omega}^{(2)}_{11{\bf p}{\bf q}}$ can be expressed approximately as,
$\bar{\omega}^{(1)}_{11{\bf p}{\bf q}}$ $=$ $\bar{\omega}_{1{\bf q}+{\bf p}}+\bar{\omega}_{1{\bf q}}\approx b_{\bf q}p^{2}+2\bar{\omega}_{1{\bf q}},$ (35a)
$\bar{\omega}^{(2)}_{11{\bf p}{\bf q}}$ $=$ $\bar{\omega}_{1{\bf q}+{\bf p}}-\bar{\omega}_{1{\bf q}}\approx b_{\bf q}p^{2},$ (35b)
with $b_{\bf q}=d^{2}\bar{\omega}_{1{\bf q}}/(d^{2}q)$. The above results in Eq. (35) show that as in the case of the the overdoped strange-metal phaseMa23 , the full effective spin propagator $P({\bf k},{\bf p},{\bf k}^{\prime},\omega)$ in Eq. (34) scales with $p^{2}$, and then when the electron umklapp scattering kicks in, the energy scale is proportional to $\Delta^{2}_{p}$ due to the presence of this $p^{2}$ scaling in the full effective spin propagator (34). In this case, $T_{\rm scale}=\bar{b}\Delta^{2}_{p}$ can be identified as the low-temperature scale, with the average value $\bar{b}=(1/N)\sum\limits_{{\bf q}\in\{{\bf k}_{\rm A}\}}b({\bf q})$ that is a constant at a given doping, where the summation ${\bf q}\in\{{\bf k}_{\rm A}\}$ is restricted to the extremely small area $\{{\bf k}_{\rm A}\}$ around the ${\bf k}_{\rm A}$ point of BZ. However, this low-temperature scale $T_{\rm scale}$ is strongly doping dependent. To see this doping dependence of $T_{\rm scale}$ more clearly, we plot $T_{\rm scale}$ as a function of doping in Fig. 9, where $T_{\rm scale}$ is relatively high at around the slightly underdoped region, then it decreases when the doping is increased in the heavily underdoped region, and is reduced to a very low temperature in the overdoped regime. Moreover, $T_{\rm scale}$ as a function of doping in Fig. 9 in the underdoped regime presents a similar behavior of the antinodal spin pseudogap crossover temperature $T^{*}_{\rm s}$ shown in Fig. 2b, suggesting a possible correlation between $T_{\rm scale}$ and the opening of the antinodal spin pseudogap below $T^{*}_{\rm s}$ in the underdoped pseudogap phase.
With the help of the above full effective spin propagator $P({\bf k},{\bf p},{\bf k}^{\prime},\omega)$ in Eq. (34) and the doping dependence of $T_{\rm scale}$ in Fig. 9, we now follow the similar analysis carried out in the overdoped strange-metal phaseMa23 to show that two primary regions of the low-temperature resistivity need to be distinguished:
(i) in the low-temperature T-quadratic region ($T<T_{\rm scale}$) in the underdoped regime, the kernel function $F(\theta,\theta^{\prime})$ is reduced as $F(\theta,\theta^{\prime})\propto T^{2}$, which naturally produces a low-temperature T-quadratic resistivity $\rho(T)\propto T^{2}$ as shown in Fig. 5;
(ii) in the low-temperature T-linear region ($T>T_{\rm scale}$) in the overdoped regime, the kernel function $F(\theta,\theta^{\prime})$ is reduced as $F(\theta,\theta^{\prime})\propto T$, which naturally induces a T-linear resistivity $\rho(T)\propto T$ as we have shown in Ref. Ma23, . In particular, this low-temperature scale $T_{\rm scale}$ in the overdoped regime is very low Ma23 due to the absence of the antinodal spin pseudogap as shown in Fig. 2.
IV Summary and discussion
Within the framework of the kinetic-energy-driven superconductivity, we have rederived the full spin propagator, the full charge-carrier propagator, and the full electron propagator in the normal-state of cuprate superconductors, where the spin self-energy (then the spin pseudogap) is obtained explicitly in terms of the collective charge-carrier mode in the particle-hole channel, and the charge-carrier normal self-energy (then the charge-carrier pseudogap) is obtained explicitly in terms of the spin excitation mode, while the electron normal self-energy (then the normal-state pseudogap) originated from the charge-carrier self-energy (then the charge-carrier pseudogap) is due to the charge-spin recombination. Moreover, we have also shown that (i) the spin excitation energy dispersion is anisotropically renormalized due to the momentum dependence of the spin pseudogap. In particular, the antinodal spin pseudogap effect is particularly notable in the slightly underdoped region, and then this antinodal spin pseudogap rapidly decreases with the increase of doping in the heavily underdoped region, eventually abrupt disappearing at around the optimal doping; (ii) the electronic density of state at around the antinodal region is gapped out by the normal-state pseudogap, and then the closed EFS contour is truncated to a set of four disconnected Fermi arcs centered at around the nodal region. By virtue of this full spin propagator and the reconstructed EFS, we have investigated the low-temperature electrical transport in the underdoped pseudogap phase of cuprate superconductors, where the scattering rate originated from the umklapp scattering between electrons by the exchange of the full effective spin propagator is derived within the framework of the Boltzmann transport theory. Our results show that the dominant contribution to the low-temperature resistivity mainly comes from the antinodal umklapp scattering. In particular, a low temperature $T_{\rm scale}$ scales with $\Delta^{2}_{p}$ in the underdoped regime due to the opening of the momentum dependence of the spin pseudogap. Moreover, this $T_{\rm scale}$ as a function of doping presents a similar behavior of the antinodal spin pseudogap crossover temperature $T^{*}_{\rm s}$, i.e., $T_{\rm scale}$ decreases with the increase of doping in the underdoped regime, and then is reduced to a very low temperature in the overdoped regime, which suggests a possible correlation between $T_{\rm scale}$ and the opening of the antinodal spin pseudogap below $T^{*}_{\rm s}$. In the underdoped regime, the resistivity exhibts a T-quadratic behaviour in the low-temperature region below $T_{\rm scale}$, where the strength of the T-quadratic resistivity decreases with the increase of doping. However, in the overdoped regime, the resistivity is T-linear in the low-temperature region above $T_{\rm scale}$. The current results together with the recent resultsMa23 of the low-temperature T-linear resistivity in the overdoped regime therefore show that (i) the electron umklapp scattering from a spin excitation associated with the antinodes leads to the T-linear resistivity in the weak coupling overdoped regime; (ii) as this electron umklapp scattering flows to the strong coupling underdoped regime, the opening of the momentum dependence of the spin pseudogap lowers the density of states of the spin excitations at around the antinodal region in the response to the intense umklapp scattering, which reduces the strength of the umklapp scattering from electronic states into the antinodal region, and therefore leads to the low-temperature T-quadratic form of the umklapp scattering rate. Concomitantly, the low-temperature resistivity exhibits a dramatic switch from the T-linear behaviour in the overdoped strange-metal phase to the T-quadratic behaviour in the underdoped pseudogap phase.
It should be emphasized that the full effective spin propagator $P({\bf k},{\bf p},{\bf k}^{\prime},\omega)$ in Eq. (24) is obtained in the underdoped regime, where the AFLRO correlation is absent, although the AFSRO correlation survives from the underdoped regime to the overdoped regime Fujita12 as we have mentioned in subsection II.4. In this case, the umklapp scattering between electrons in Eq. (18) by the exchange of this full effective spin propagator is better suited for the description of the electrical transport in the underdoped regime. However, in the extremely light-doped regime ($\delta\leq 0.05$), the AFLRO correlation has been identified by INS, muon spin rotation, and other measurements Kastner98 ; Fujita12 ; Vaknin87 ; Brewer88 . Moreover, the strength of the AFLRO correlation as a function of the doping concentration has been established by NMR measurementsKitaoka88 , where AFLRO is destructed by a few percent of the doping concentration ($\sim 0.05$). In this extremely light-doped regime, the electronic state of the system with some special features appear, in particular, the low-temperature resistivity shows a upturn Ando01 ; Ando04a ; Ando04 . Although the mechanism causing this resistivity upturn remains unclear, it is possible that this striking resistivity upturn can be also described in terms of the umklapp scattering between electrons by the exchange of the effective spin propagator, where the spin propagator should give a suitable description of the special magnetic properties with the AFLRO correlation. These and the related issues are under investigation now.
Acknowledgements
This work is supported by the National Key Research and Development Program of China under Grant Nos. 2023YFA1406500 and 2021YFA1401803, and the National Natural Science Foundation of China (NSFC) under Grant Nos. 12274036 and 12247116. H.G. acknowledge support from NSFC grant Nos. 11774019 and 12074022.
Appendix A Derivation of full charge-carrier, full spin, and full electron propagators
In this Appendix, the main goal is to derive the full charge-carrier propagator $g({\bf k},\omega)$ in Eq. (3), the full spin propagator $D({\bf k},\omega)$ in Eq. (5), and the full electron propagator $G({\bf k},\omega)$ in Eq. (11) of the main text. Following the fermion-spin transformation (2), the $t$-$J$ model (1) can be rewritten as,
$H$ $=$ $\sum_{\langle l\hat{\eta}\rangle}t(h^{\dagger}_{l+\hat{\eta}\uparrow}h_{l\uparrow}S^{+}_{l}S^{-}_{l+\hat{\eta}}+h^{\dagger}_{l+\hat{\eta}\downarrow}h_{l\downarrow}S^{-}_{l}S^{+}_{l+\hat{\eta}})$ (36)
-- $\sum_{\langle l\hat{\tau}\rangle}t^{\prime}(h^{\dagger}_{l+\hat{\tau}\uparrow}h_{l\uparrow}S^{+}_{l}S^{-}_{l+\hat{\tau}}+h^{\dagger}_{l+\hat{\tau}\downarrow}h_{l\downarrow}S^{-}_{l}S^{+}_{l+\hat{\tau}})$
-- $\mu_{\rm h}\sum_{l\sigma}h^{\dagger}_{l\sigma}h_{l\sigma}+J_{\rm eff}\sum_{\langle l\hat{\eta}\rangle}{\bf S}_{l}\cdot{\bf S}_{l+\hat{\eta}},$
with $J_{{\rm eff}}=(1-\delta)^{2}J$, the charge-carrier doping concentration $\delta=\langle h^{\dagger}_{l\sigma}h_{l\sigma}\rangle=\langle h^{\dagger}_{l}h_{l}\rangle$, and the charge-carrier chemical potential $\mu_{\rm h}$. The above $t$-$J$ model (36) therefore describes a doped AF insulator as a sparse density of the charge carriers moving in a background of an AF coupled square lattice of spins, while the motion of the electrons rearranges the spin configuration leading to the strong coupling between the charge and spin degrees of freedom of the constrained electron.
A.1 Full charge-carrier propagator
In the early studiesFeng15 ; Feng0306 ; Feng12 ; Feng15a , it has been shown that the interaction between the charge carriers directly from the kinetic energy of the $t$-$J$ model (36) by the exchange of the spin excitation generates the charge-carrier pairing state in the particle-particle channel. According to these early studies, the self-consistent equations that are satisfied by the full charge-carrier diagonal and off-diagonal propagators in the charge-carrier pairing state have been evaluated in terms of the Eliashberg formalismEliashberg60 ; Scalapino66 , and then in the charge-carrier normal-state, these self-consistent equations are reduced asFeng15 ; Feng12 ,
$g({\bf k},\omega)=g^{(0)}({\bf k},\omega)+g^{(0)}({\bf k},\omega)\Sigma^{({\rm h})}_{\rm ph}({\bf k},\omega)g({\bf k},\omega),~{}~{}~{}~{}$ (37)
with the charge-carrier propagator of the $t$-$J$ model (36) in the MF approximation $g^{(0)-1}({\bf k},\omega)=\omega-\xi_{\bf k}$. From the above self-consistent equation (37), the full charge-carrier propagator can be expressed explicitly as,
$g({\bf k},\omega)={1\over\omega-\xi_{\bf k}-\Sigma^{({\rm h})}_{\rm ph}({\bf k},\omega)},$ (38)
which is the same as quoted in Eq. (3) of the main text. Moreover, the charge-carrier normal self-energy $\Sigma^{({\rm h})}_{\rm ph}({\bf k},\omega)$ has been derived as Feng15 ; Feng12 ,
$\Sigma^{({\rm h})}_{\rm ph}({\bf k},i\omega_{n})$ $=$ ${1\over N}\sum_{\bf p}{1\over\beta}\sum_{ip_{m}}g({\bf p}+{\bf k},ip_{m}+i\omega_{n})$ (39)
$\times$ $P^{(0)}({\bf k},{\bf p},ip_{m}),~{}~{}~{}$
with the number of lattice sites N, the fermion and bosonic Matsubara frequencies $\omega_{n}$ and $p_{m}$, respectively, and the MF effective spin propagator,
$P^{(0)}({\bf k},{\bf p},\omega)={1\over N}\sum_{\bf q}\Lambda^{2}_{{\bf p}+{\bf q}+{\bf k}}\Pi({\bf p},{\bf q},\omega),$ (40)
where the MF spin bubble $\Pi({\bf p},{\bf q},\omega)$ is a convolution of two MF spin propagators, and can be expressed as,
$\Pi({\bf p},{\bf q},ip_{m})={1\over\beta}\sum_{iq_{m}}D^{(0)}({\bf q},iq_{m})D^{(0)}({\bf q}+{\bf p},iq_{m}+ip_{m}),~{}~{}~{}$ (41)
with the bosonic Matsubara frequency $q_{m}$, and the MF spin propagator,
$D^{(0)}({\bf k},\omega)={B_{\bf k}\over\omega^{2}-\omega^{2}_{\bf k}}={B_{\bf k}\over 2\omega_{\bf k}}\left({1\over\omega-\omega_{\bf k}}-{1\over\omega+\omega_{\bf k}}\right).~{}~{}~{}~{}~{}$ (42)
With the help of the above spin propagator (42), the charge-carrier normal self-energy $\Sigma^{({\rm h})}_{\rm ph}({\bf k},\omega)$ in Eq. (39) can be obtained asFeng15 ; Feng12 ,
$\Sigma^{({\rm h})}_{\rm ph}({\bf k},\omega)$ $=$ ${1\over N^{2}}\sum_{{\bf pp}^{\prime}\mu\nu}(-1)^{\nu+1}Z^{({\rm h})}_{\rm F}\Omega_{{\bf p}{\bf p}^{\prime}{\bf k}}$ (43)
$\times$ ${F^{({\rm h})}_{\mu\nu}({\bf p},{\bf p}^{\prime},{\bf k})\over\omega+(-1)^{\mu+1}\omega^{(\nu)}_{{\bf p}{\bf p}^{\prime}}-\bar{\xi}_{{\bf p}+{\bf k}}},~{}~{}~{}$
with $\mu~{}(\nu)=1,~{}2$, $\omega^{(\nu)}_{{\bf p}{\bf p}^{\prime}}=\omega_{{\bf p}+{\bf p}^{\prime}}-(-1)^{\nu}\omega_{\bf p^{\prime}}$, $\Omega_{{\bf p}{\bf p}^{\prime}{\bf k}}=\Lambda^{2}_{{\bf p}+{\bf p}^{\prime}+{\bf k}}B_{{\bf p}^{\prime}}B_{{\bf p}+{\bf p}^{\prime}}/[4\omega_{{\bf p}^{\prime}}\omega_{{\bf p}+{\bf p}^{\prime}}]$, $F^{({\rm h})}_{\mu\nu}({\bf p},{\bf p}^{\prime},{\bf k})=n_{\rm F}[(-1)^{\mu+1}\bar{\xi}_{{\bf p}+{\bf k}}]n^{(\nu)}_{{\rm 1B}{{\bf p}{\bf p}^{\prime}}}+n^{(\nu)}_{{\rm 2B}{{\bf p}{\bf p}^{\prime}}}$, $n^{(\nu)}_{{\rm 1B}{{\bf p}{\bf p}^{\prime}}}=1+n_{\rm B}(\omega_{{\bf p}^{\prime}+{\bf p}})+n_{\rm B}[(-1)^{\nu+1}\omega_{\bf p^{\prime}}]$, and $n^{(\nu)}_{{\rm 2B}{{\bf p}{\bf p}^{\prime}}}=n_{\rm B}(\omega_{{\bf p}^{\prime}+{\bf p}})n_{\rm B}[(-1)^{\nu+1}\omega_{\bf p^{\prime}}]$.
A.2 Full spin propagator
Starting from the $t$-$J$ model (36) in the fermion-spin representation, the full spin propagator in the normal-state has been evaluated as Kuang15 ; Yuan01 ; Feng98 ,
$D({\bf k},\omega)$ $=$ ${1\over D^{(0)-1}({\bf k},\omega)-\Sigma^{({\rm s})}_{\rm ph}({\bf k},\omega)},~{}~{}~{}~{}~{}$ (44)
where the spin self-energy in the normal-state is derived in terms of the collective charge-carrier mode in the particle-hole channel as,
$\Sigma^{({\rm s})}_{\rm ph}({\bf k},\omega)$ $=$ $-{1\over N^{2}}\sum_{\bf pq}\Omega^{({\rm s})}_{{\bf k}{\bf p}{\bf q}}{F^{\rm(s)}({\bf k},{\bf p},{\bf q})\over\omega^{2}-[\omega_{{\bf q}+{\bf k}}-(\bar{\xi}_{{\bf p}+{\bf q}}-\bar{\xi}_{{\bf p}})]^{2}},$
with the vertex function $\Omega^{({\rm s})}_{{\bf k}{\bf p}{\bf q}}$ and the function $F^{\rm(s)}({\bf k},{\bf p},{\bf q})$ that are given by,
$\Omega^{({\rm s})}_{{\bf k}{\bf p}{\bf q}}$ $=$ ${B_{{\bf q}+{\bf k}}\over\omega_{{\bf q}+{\bf k}}}[Z^{({\rm h})}_{\rm F}]^{2}(\Lambda^{2}_{{\bf k}-{\bf p}}+\Lambda^{2}_{{\bf p}+{\bf q}+{\bf k}}),~{}~{}~{}~{}~{}~{}$ (46a)
$F^{\rm(s)}({\bf k},{\bf p},{\bf q})$ $=$ $[\omega_{{\bf q}+{\bf k}}-(\bar{\xi}_{{\bf p}+{\bf q}}-\bar{\xi}_{{\bf p}})]\{n_{\rm B}(\omega_{{\bf q}+{\bf k}})[n_{\rm F}(\bar{\xi}_{{\bf p}})$ (46b)
-- $n_{\rm F}(\bar{\xi}_{{\bf p}+{\bf q}})]-[1-n_{\rm F}(\bar{\xi}_{{\bf p}})]n_{\rm F}(\bar{\xi}_{{\bf p}+{\bf q}})\},~{}~{}~{}~{}~{}~{}~{}~{}$
respectively. Substituting the above spin self-energy in Eq. (A.2) into Eq. (44), the full spin propagator in the normal-state can be expressed as,
$D({\bf k},\omega)$ $=$ ${B_{\bf k}\over\omega^{2}-\omega^{2}_{\bf k}-B_{\bf k}\Sigma^{({\rm s})}_{\rm ph}({\bf k},\omega)},~{}~{}~{}~{}~{}$ (47)
which is the same as quoted in Eq. (5) of the main text.
A.3 Full electron propagator
In order to obtain the electron propagator, the full charge-spin recombination scheme has been developedFeng15a , where it has been shown that the coupling form between the electron and spin excitation in the $t$-$J$ model in the normal-state is the same as that between the charge carrier and spin excitation in Eq. (37), and then the self-consistent equation satisfied by the full electron propagator in the normal-state can be derived directly asFeng15a ,
$G({\bf k},\omega)=G^{(0)}({\bf k},\omega)+G^{(0)}({\bf k},\omega)\Sigma_{\rm ph}({\bf k},\omega)G({\bf k},\omega),~{}~{}~{}~{}$ (48)
where $G^{(0)-1}({\bf k},\omega)=\omega-\varepsilon_{\bf k}$ is the electron propagator of the $t$-$J$ model (1) in the tight-binding approximation. From the above equation (48), the full electron propagator $G({\bf k},\omega)$ can be expressed explicitly as,
$G({\bf k},\omega)={1\over\omega-\varepsilon_{\bf k}-\Sigma_{\rm ph}({\bf k},\omega)},$ (49)
which is the same as quoted in Eq. (11) of the main text. In particular, the electron normal self-energy,
$\Sigma_{\rm ph}({\bf k},i\omega_{n})$ $=$ ${1\over N}\sum_{\bf p}{1\over\beta}\sum_{ip_{m}}G({\bf p}+{\bf k},ip_{m}+i\omega_{n})$ (50)
$\times$ $P^{(0)}({\bf k},{\bf p},ip_{m}),~{}~{}~{}$
has been evaluated asFeng15a ,
$\Sigma_{\rm ph}({\bf k},\omega)$ $=$ ${1\over N^{2}}\sum_{{\bf pp^{\prime}}\mu\nu}(-1)^{\nu+1}Z_{\rm F}\Omega_{\bf pp^{\prime}k}$ (51)
$\times$ ${F_{\mu\nu}({\bf p},{\bf p}^{\prime}{\bf k})\over\omega+(-1)^{\mu+1}\omega_{\nu{\bf p}{\bf p}^{\prime}}-\bar{\varepsilon}_{{\bf p}+{\bf k}}},~{}~{}~{}$
with the function,
$F_{\mu\nu}({\bf p},{\bf p}^{\prime}{\bf k})=n_{\rm F}[(-1)^{\mu+1}\bar{\varepsilon}_{{\bf p}+{\bf k}}]n^{(\nu)}_{{\rm 1B}{\bf pp^{\prime}}}+n^{(\nu)}_{{\rm 2B}{\bf pp^{\prime}}},~{}~{}~{}~{}~{}$ (52)
where $n^{(\nu)}_{{\rm 1B}{\bf pp^{\prime}}}=1+n_{\rm B}(\omega_{{\bf p}^{\prime}+{\bf p}})+n_{\rm B}[(-1)^{\nu+1}\omega_{\bf p^{\prime}}]$, and $n^{(\nu)}_{{\rm 2B}{\bf pp^{\prime}}}=n_{\rm B}(\omega_{{\bf p}^{\prime}+{\bf p}})n_{\rm B}[(-1)^{\nu+1}\omega_{\bf p^{\prime}}]$. This electron normal self-energy characterizes a competition between the kinetic energy and magnetic energy in the $t$-$J$ model (1).